### v1

# Modular resurgent structures

Abstract
The theory of resurgence uniquely associates a factorially divergent formal power series with a collection of non-perturbative, exponential-type corrections paired with a set of complex numbers, known as Stokes constants. When the Borel plane displays a single infinite tower of singularities, the secondary resurgent series are trivial, and the Stokes constants are coefficients of an $L$-function, a rich analytic number-theoretic fabric underlies the resurgent structure of the asymptotic series. We propose a new paradigm of modular resurgence that focuses on the role of the Stokes constants and the interplay of the $q$-series acting as their generating functions with the corresponding $L$-functions. Guided by two pivotal examples arising from topological string theory and the theory of Maass cusp forms, we introduce the notion of modular resurgent series, which we conjecture to have specific summability properties as well as to be closely related to quantum modular forms.
1 Introduction
Resurgence, summability, and modularity are very active domains of research involving different communities, which have so far mostly focused on the study of various examples mainly coming from quantum Chern–Simons theory and the quantum invariants of knots and 3-manifolds [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. There is no general theory that integrates these three research areas into a comprehensive framework. In this paper, we aim to address this gap.
We take advantage of the abundant amount of information gathered from the in-depth analysis of the resurgence of the spectral trace of the toric Calabi–Yau (CY) threefold known as local ${\mathbb{P}}^{2}$ performed by the second author in [18]. This satisfies a full-fledged strong-weak resurgent symmetry and special modularity and summability properties that we present in the companion paper [19]. Generalizing the cardinal features of this example, we give a characterization of those resurgent asymptotic series that produce holomorphic quantum modular forms and, at the same time, whose median resummation is effective. This newly proposed paradigm goes under the name of modular resurgence and deeply relies on the fundamental role played by Stokes constants and the $q$-series acting as their generating functions. Our results naturally lead us to present the definition of a modular resurgent series (see Definitions 3.1 and 3.2), whose Borel plane displays a single infinite tower of singularities, the secondary resurgent series are trivial, and the Stokes constants are coefficients of an $L$-function. The resurgent structure of such a series is called a modular resurgent structure.
We show that a global network of exact relations connects the full resurgent structures of pairs of modular resurgent series, which is built upon the analytic properties of the $L$-functions whose coefficients are the Stokes constants. Guided by a diverse and conspicuous set of examples that provide supporting evidence, we conjecture that those $q$-series whose perturbative series is modular resurgent are holomorphic quantum modular forms and are reconstructed by the median resummation of their asymptotic expansions (see Conjectures 1 and 2). Beyond the complete, original example of the spectral trace of local ${\mathbb{P}}^{2}$ of [19], our statements are proven for a large class of modular resurgent series that originate from the theory of Maass cusp forms.
Although conjectural, our paradigm of modular resurgence is supported by evidence coming from the study of quantum invariants of knots and 3-manifolds, combinatorics, and topological string theory on toric CY threefolds.
  • In the study of quantum invariants of knots, the example of the trefoil knot and other families of toric knots have been discussed from the resurgence, summability, and modularity perspectives separately. Evidence of our conjectures is present in the following works [1, 20]. Other pieces of evidence come from the study of 3-manifold invariants, such as the Witten–Reshetikhin–Turaev (WRT) invariants. The authors of [2] prove that the Ohtsuki series $Z_{0}$ for Seifert fibered homology spheres satisfies the conditions of what we define here as a modular resurgent series. Its median resummation reconstructs the normalized WRT invariants $\tilde{Z}_{0}$, which are related to the Gukov–Pei–Putrov–Vafa (GPPV) invariants $\hat{Z}_{0}$ of [21]. Then, the results of [13, 14] show that the GPPV invariants of certain plumbed 3-manifolds are quantum modular forms.
  • From combinatorics, families of $q$-hypergeometric series [22] provide examples of quantum modular forms, among which the example of $\sigma,\sigma^{*}$ first discussed by Zagier in [6, Example 1]. The resurgent structure of $\sigma,\sigma^{*}$, presented in [23], has been studied by the first author and Kontsevich and is of modular type. Furthermore, we prove here that Maass cusp forms with spectral parameter $1/2$ provide a complete model of modular resurgence.
  • In topological string theory, and particularly in the context of the TS/ST correspondence, the only example we are aware of comes from the results obtained in [18, 19] on the resurgence of the spectral trace of local ${\mathbb{P}}^{2}$ at strong and weak coupling, although we expect the existence of an appropriate generalization to other toric CY threefolds and higher-order fermionic spectral traces.
This paper is organized as follows. In Section 2, we review the basic ingredients in the resurgent analysis of formal power series and the definitions of (holomorphic) quantum modular forms. In Section 3, we delve into the analytic number-theoretic characterization of those resurgent structures that give rise to quantum modular forms according to the newly introduced paradigm of modular resurgence. We present the definition of a modular resurgent series, which we conjecture to have peculiar modularity and summability properties. In Section 4, we review the example of the spectral trace of local ${\mathbb{P}}^{2}$, briefly summarizing the strong-weak resurgent symmetry of [19], and prove our conjectures for a set of examples originating from the theory of Maass cusp forms. Finally, in Section 5, we conclude and mention further perspectives to be addressed in future works.
2 Basics of resurgence and modularity
In this section, we review the necessary background notions in resurgence and modularity. The resurgent analysis of factorially divergent formal power series [24] unveils a universal mathematical structure of non-perturbative sectors involving a collection of numerical data called Stokes constants. See [25, 26, 27] for a formal introduction to the theory of resurgence. Classical modular forms have a long history and are of interest in different fields. Generally speaking, they outline the invariant properties of certain $q$-series under the action of the modular group. However, there are examples of $q$-series that are close to being modular forms and whose failure of modularity can be described explicitly. These belong to the class of quantum modular forms introduced by Zagier in [6]. We refer to [28, Chapter IV] and references therein for an overview.
2.1 Resurgent asymptotic series
Let $z$ be a formal variable. The Borel transform is a map acting on $z$-monomials as
$\mathcal{B}[z^{n-\alpha}]:=\frac{\zeta^{n-\alpha-1}}{\Gamma(n-\alpha)}\,,\quad n\in{\mathbb{Z}}_{\geq 0}\,,\quad\alpha\in{\mathbb{R}}\backslash{\mathbb{Z}}_{\geq 0}\,,$ (1)
where $\zeta$ is a new formal variable and $\Gamma(n-\alpha)$ is the gamma function. The action of the Borel transform above is extended by linearity to all formal power series in $z^{-\alpha}{\mathbb{C}}[\![z]\!]$, while the Borel transform of the identity is conventionally denoted with the symbol $\delta:=\mathcal{B}[1]$. Let $\phi(z)$ be a Gevrey-1 asymptotic series of the form
$\phi(z)=z^{-\alpha}\sum_{n=0}^{\infty}a_{n}z^{n}\in z^{-\alpha}{\mathbb{C}}[\![z]\!]\,,\quad|a_{n}|\leq{\cal A}^{-n}n!\quad n\gg 1\,,$ (2)
where $\alpha\in{\mathbb{R}}\backslash{\mathbb{Z}}_{\geq 0}$ and ${\cal A}\in{\mathbb{R}}_{>0}$. Its Borel transform, which we denote for simplicity by $\hat{\phi}(\zeta):=\mathcal{B}\phi(\zeta)$, can be written as
$\hat{\phi}(\zeta)=\sum_{k=0}^{\infty}\frac{a_{k}}{\Gamma(k-\alpha)}\zeta^{k-\alpha-1}\in\zeta^{-\alpha}{\mathbb{C}}\{\zeta\}\,,$ (3)
which is holomorphic in an open neighborhood of $\zeta=0$ of radius $|{\cal A}|$. When extended to the complex $\zeta$-plane, known as Borel plane, $\hat{\phi}(\zeta)$ shows a (possibly infinite) set of singularities $\zeta_{\omega}\in{\mathbb{C}}$, which we label by the index $\omega\in\Omega$. A ray in the Borel plane of the form
${\cal C}_{\theta_{\omega}}={\rm e}^{\text{i}\theta_{\omega}}{\mathbb{R}}_{+}\,,\quad\theta_{\omega}=\arg(\zeta_{\omega})\,,$ (4)
which starts at the origin and passes through a singularity $\zeta_{\omega}$, is called a Stokes ray. The Borel plane is partitioned into sectors bounded by the Stokes rays in such a way that the Borel transform converges to a generally different holomorphic function in each of these sectors.
We recall that the Gevrey-1 series $\phi(z)$ is called resurgent if its Borel transform $\hat{\phi}(\zeta)$ can be endlessly analytically continued. Namely, for every $L>0$, there is a finite set of points $\Omega_{L}$ in the Riemann surface of $\zeta^{\alpha}$ such that $\hat{\phi}(\zeta)$ can be analytically continued along any path that avoids $\Omega_{L}$ and has length at most $L$. If, additionally, its Borel transform has only simple poles and logarithmic branch points, then it is called simple resurgent. Let us assume that the formal power series $\phi(z)$ in Eq. (2) is simple resurgent. If the singularity $\zeta_{\omega}$ is a simple pole, the local expansion of the Borel transform in Eq. (3) around it has the form
$\hat{\phi}(\zeta)=-\frac{S_{\omega}}{2\pi\text{i}(\zeta-\zeta_{\omega})}+\text{regular in $\zeta-\zeta_{\omega}$}\,,$ (5)
where $S_{\omega}\in{\mathbb{C}}$ is the Stokes constant at $\zeta_{\omega}$. Whereas, if the singularity $\zeta_{\omega}$ is a logarithmic branch point, the local expansion of the Borel transform in Eq. (3) around it has the form
$\hat{\phi}(\zeta)=-\frac{S_{\omega}}{2\pi\text{i}}\log(\zeta-\zeta_{\omega})\hat{\phi}_{\omega}(\zeta-\zeta_{\omega})+\text{regular in $\zeta-\zeta_{\omega}$}\,,$ (6)
where again $S_{\omega}\in{\mathbb{C}}$ is the corresponding Stokes constant. If we introduce the variable $\xi=\zeta-\zeta_{\omega}$, the function
$\hat{\phi}_{\omega}(\xi)=\sum_{k=0}^{\infty}\hat{a}_{k,\omega}\xi^{k-\beta}\in\xi^{-\beta}{\mathbb{C}}\{\xi\}\,,$ (7)
where $\subscript\beta\in{\mathbb{R}}\backslash{\mathbb{Z}}_{+}$, is locally analytic at $\xi=0$ and can be regarded as the Borel transform of the Gevrey-1 asymptotic series
$\phi_{\omega}(z)=z^{-\beta}\sum_{n=0}^{\infty}a_{n,\omega}z^{n}\in z^{-\beta}{\mathbb{C}}[\![z]\!],\,\quad a_{n,\omega}=\Gamma(n-\beta+1)\,\hat{a}_{n,\omega}\,.$ (8)
Note that the value of the Stokes constant $S_{\omega}$ depends on the normalization of $\phi_{\omega}(z)$.
If the analytic continuation of the Borel transform $\hat{\phi}(\zeta)$ in Eq. (3) does not grow too fast at infinity [^1] Roughly, we require that the Borel transform grows at most exponentially in an open sector of the Borel plane containing the angle $\theta$., its Laplace transform at an arbitrary angle $\theta$ in the Borel plane gives the Borel–Laplace sum of the original, divergent formal power series $\phi(z)$, denoted by $s_{\theta}(\phi)(z)$. Explicitly,
[…]
whose asymptotics near the origin reconstructs $\phi(z)$. If the Borel–Laplace sum in Eq. (9) exists in some region of the complex $z$-plane, we say that the series $\phi(z)$ is Borel–Laplace summable along the direction $\theta$. Note that the Borel–Laplace sum inherits the sectorial structure of the Borel transform. It is a locally analytic function with discontinuities across the special rays identified by
$\arg(z)=\arg(\zeta_{\omega})\,,\quad\omega\in\Omega\,.$ (10)
The discontinuity across an arbitrary ray ${\cal C}_{\theta}={\rm e}^{\text{i}\theta}{\mathbb{R}}_{+}$ in the complex $z$-plane is the difference between the Borel–Laplace sums along two rays in the complex $\zeta$-plane that lie slightly above and slightly below ${\cal C}_{\theta}$. Namely,
[…]
where $\theta_{\pm}=\theta\pm\epsilon$ for some small positive angle $\epsilon$. A standard contour deformation argument shows that the two lateral Borel–Laplace sums differ by exponentially small terms. More precisely, if the Borel transform $\hat{\phi}(z)$ has only simple poles, then we have that
$\mathrm{disc}_{\theta}\phi(z)=\sum_{\omega\in\Omega_{\theta}}S_{\omega}{\rm e}^{-\zeta_{\omega}/z}\,,$ (12)
where the index $\omega\in\Omega_{\theta}$ labels the singularities $\zeta_{\omega}$ such that $\arg(\zeta_{\omega})=\theta$ and the complex numbers $S_{\omega}$ are the same Stokes constants that appear in Eq. (5). Similarly, if there are only logarithmic branch points, we have that
$\mathrm{disc}_{\theta}\phi(z)=\sum_{\omega\in\Omega_{\theta}}S_{\omega}{\rm e}^{-\zeta_{\omega}/z}s_{\theta_{-}}(\phi_{\omega})(z)\,,$ (13)
where again the complex numbers $S_{\omega}$ are the Stokes constants appearing in Eq. (6), while $\phi_{\omega}(z)$ is the formal power series in Eq. (8).
In the study of the summability properties of factorially divergent formal power series that arise as asymptotic limits of well-defined holomorphic functions, it is natural to ask the question of which summability method is efficient in giving back the original function. In quantum Chern–Simons theory, Costin and Garoufalidis conjectured that the asymptotic expansions at roots of unity of quantum invariants, such as the WRT invariant for knots and the Kashaev invariant for $3$-manifolds, are Gevrey-1 series whose median resummation reconstructs the non-perturbative invariants [1, Conjecture 1]. As we show in [19], more evidence of the efficacy of the median resummation comes from the study of the spectral traces of the quantum operators obtained by quantizing the mirror curve to a toric CY threefold. We recall that the median resummation of the Gevrey-1 asymptotic series $\phi(z)$ in Eq. (2) across an arbitrary ray ${\cal C}_{\theta}$ in the complex $z$-plane is the average of the two lateral Borel–Laplace sums $s_{\theta_{\pm}}(\phi)(z)$, that is,
$\mathcal{S}^{\text{med}}_{\theta}\phi(z)=\frac{s_{\theta_{+}}(\phi)(z)+s_{\theta_{-}}(\phi)(z)}{2}\,,$ (14)
which is an analytic function for $\arg z\in(\theta-\frac{\pi}{2},\theta+\frac{\pi}{2})$. Equivalently, we can write
$\mathcal{S}^{\text{med}}_{\theta}\phi(z)=\begin{cases}s_{\theta_{-}}(\phi)(z)+\frac{1}{2}\,\mathrm{disc}_{\theta}\phi(z)\,,\quad&\Re\left({\rm e}^{-{\rm i}\theta_{-}}z\right)>0\,,\\ &\\ s_{\theta_{+}}(\phi)(z)-\frac{1}{2}\,\mathrm{disc}_{\theta}\phi(z)\,,\quad&\Re\left({\rm e}^{-{\rm i}\theta_{+}}z\right)>0\,,\end{cases}$ (15)
where $\Re\left({\rm e}^{-{\rm i}\theta}z\right)>0$. The expression above highlights how the discontinuities can be suitably interpreted as corrections to the standard Borel–Laplace transform. In addition, by varying the contour of integration of the Laplace transform, the domain of analyticity of the median resummation can be extended, which is helpful in some applications.
2.2 Quantum modular forms
A first description of quantum modular forms was given by Zagier in [6]. Namely, a function $f\colon{\mathbb{Q}}\to{\mathbb{C}}$ is called a weight-$\omega$ quantum modular form with respect to a subgroup $\Gamma\subseteq\mathsf{SL}_{2}({\mathbb{Z}})$, where $\omega$ is a fixed integer or half-integer, if the cocycle
$h_{\gamma}[f](y):=(cy+d)^{-\omega}f\left(\frac{ay+b}{cy+d}\right)-f(y)$ (16)
is better behaved than $f(y)$ for all choices of $\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma$. Conceptually, being better behaved means having better analyticity properties than the function $f(y)$ itself – e.g., being a real analytic function over ${\mathbb{R}}\setminus\{\gamma^{-1}(\infty)\}$. Quantum modular forms of weight zero are called quantum modular functions.
Some examples of quantum modular forms are built from classical modular forms. For instance, it was proven in [20] that the function $g\colon{\mathbb{Q}}\to{\mathbb{C}}$ defined by
$g(y):=q^{1/24}\Big{(}1+\sum_{n=1}^{\infty}(1-q)(1-q^{2})\ldots(1-q^{n})\Big{)}\,,\quad q={\rm e}^{2\pi{\rm i}y}\,,$ (17)
which is closely related to the Kashaev invariant for the trefoil knot [1] [^2] Other examples arising from the quantum invariants of hyperbolic knots have been discussed in the literature, although their cocycles generally have a more complicated structure [7]. It is unknown whether they are related to classical modular forms., is a weight-$3/2$ quantum modular form for $\mathsf{SL}_{2}({\mathbb{Z}})$. Its cocycle is written explicitly as
[…]
where $\eta$ is the Dedekind eta function and $\gamma\in\mathsf{SL}_{2}({\mathbb{Z}})$. In particular, the cocycle $h_{\gamma}[g]$ in Eq. (18) is real analytic on ${\mathbb{R}}\setminus\{\gamma^{-1}(\infty)\}$ and is expressed in terms of the original weight-1/2 modular form $\eta$. This is a common feature of those quantum modular forms that arise from a classical counterpart and manifests in other examples. See [6, Examples 1, 3, 4] and [22, 12, 13, 14].
A function $f$ that is well-defined and analytic in the upper half of the complex plane, which we denote by ${\mathbb{H}}$, already has good analyticity properties. Yet, a version of quantum modularity is obtained by requiring that the cocycle $h_{\gamma}[f]$ in Eq. (16) is analytic in a domain larger than ${\mathbb{H}}$. The quantum modular forms that belong to this class of examples are called holomorphic quantum modular forms. The following definition was proposed by Zagier [29, min 21:45].
Definition 2.1.
An analytic function $f\colon{\mathbb{H}}\to{\mathbb{C}}$ is a weight-$\omega$ holomorphic quantum modular form for a subgroup $\Gamma\subseteq\mathsf{SL}_{2}({\mathbb{Z}})$, where $\omega$ is integer or half-integer, if the cocycle $h_{\gamma}[f]\colon{\mathbb{H}}\to{\mathbb{C}}$ in Eq. (16) extends holomorphically to ${\mathbb{C}}_{\gamma}:=\{y\in{\mathbb{C}}\colon cy+d\in{\mathbb{C}}\setminus{\mathbb{R}}_{\leq 0}\}$ [^3] Note that ${\mathbb{C}}_{\gamma}={\mathbb{C}}\setminus\left(-\infty;\,-d/c\right]$ when $c>0$ and ${\mathbb{C}}_{\gamma}={\mathbb{C}}\setminus\left[-d/c;\,+\infty\right)$ when $c<0$. for every $\gamma=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma$.
For instance, let $\Delta\colon{\mathbb{H}}\to{\mathbb{C}}$ be the modular discriminant of the Dedekind eta function $\eta$, which is given by
$\Delta(y):=\eta(y)^{24}=q\prod_{n=1}^{\infty}(1-q^{n})^{24}\,,\quad q={\rm e}^{2\pi{\rm i}y}\,.$ (19)
It was shown in [6, Example 2] that its Eichler integral $\tilde{\Delta}\colon{\mathbb{H}}\to{\mathbb{C}}$, which is given by
[…]
is a weight-10 holomorphic quantum modular form for $\mathsf{SL}_{2}({\mathbb{Z}})$, whose cocycle is known as
[…]
where again $\gamma\in\mathsf{SL}_{2}({\mathbb{Z}})$. In particular, the cocycle $h_{\gamma}[\tilde{\Delta}]$ in Eq. (21) is a polynomial of degree less than or equal to $10$ and therefore holomorphic on ${\mathbb{C}}$. This is indeed an exceptional example since, typically, the cocycle does not extend to the whole complex plane, but rather some cut plane, e.g., ${\mathbb{C}}\setminus{\mathbb{R}}_{\leq 0}$. More general examples can be constructed from the theory of Maass cusp forms and their periods, as we discuss in Section 4.2. In addition, and remarkably, as noticed in [19, Section 4.3], the resurgence of the spectral trace of local ${\mathbb{P}}^{2}$ is a new source of examples of holomorphic quantum modular forms. See the discussion in Section 4.1.
Note that an analogue of Definition 2.1 can be written for a function $f\colon{\mathbb{H}}_{-}\to{\mathbb{C}}$, where ${\mathbb{H}}_{-}$ denotes the lower half of the complex plane. Even more generally, a notion of vector-valued quantum modular forms on ${\mathbb{H}}\cup{\mathbb{Q}}\cup{\mathbb{H}}_{-}$ was proposed in [28, Definition 22].
Definition 2.2.
A function $f\colon{\mathbb{H}}\cup{\mathbb{Q}}\cup{\mathbb{H}}_{-}\rightarrow{\mathbb{C}}^{N}$, $N\in{\mathbb{Z}}_{>0}$, is a vector-valued quantum modular form for a subgroup $\Gamma\subseteq\mathsf{SL}_{2}({\mathbb{Z}})$ with an automorphic factor $\mathbf{j}\colon\Gamma\times{\mathbb{H}}\cup{\mathbb{Q}}\cup{\mathbb{H}}_{-}\to\mathsf{GL}_{N}({\mathbb{C}})$ and a QM-cocycle [^4] The notion of QM-cocyle is introduced in [28] to characterize matrix-valued holomorphic functions $\Omega$ that map $\gamma\in\Gamma\subseteq\mathsf{SL}_{2}({\mathbb{Z}})$ to $\Omega_{\gamma}\in\mathsf{GL}_{N}(\mathcal{O}({\mathbb{C}}_{\gamma}))$ and satisfy $\Omega_{\gamma\gamma^{\prime}}(y)=\Omega_{\gamma}(\gamma^{\prime}(y))\Omega_{\gamma^{\prime}}(y)$ for $\gamma,\gamma^{\prime}\in\Gamma$. $\Omega$ if it satisfies
$f(\gamma(y))=\mathbf{j}(\gamma;\,y)\Omega_{\gamma}(y)f(y)\,,$ (22)
where $\Omega_{\gamma}\in\mathsf{GL}_{N}(\mathcal{O}({\mathbb{C}}_{\gamma}))$, for every $\gamma\in\Gamma$.
One could further generalize the above definition to consider solutions to $q$-difference equations, which leads to the notion of quantum Jacobi forms [28, Definition 23]. The prominent examples are the $q$-Pochhammer symbols, as discussed in [28, Section 8.11, Theorem A-41], whose cocycle is expressed in terms of Faddeev’s quantum dilogarithm. [^5] The same feature is also observed in the example of local ${\mathbb{P}}^{2}$ [19].
3 Modular resurgence
In this section, we consider a class of Gevrey-1 asymptotic series that replicate the central features of the resurgence of the spectral trace of local ${\mathbb{P}}^{2}$, which has been extensively studied in [18, 19]. We propose a new perspective on their resurgent structures by taking the point of view of the $q$-series acting as generating functions of the Stokes constants. In particular, a cardinal role in this paradigm of modular resurgence is played by the $L$-functions whose coefficients are given by the Stokes constants and their functional equation. We describe a global exact resurgent symmetry connecting pairs of what we define as modular resurgent series and present a conjectural relation between the latter and quantum modular forms.
3.1 New perspectives on resurgence via $L$-functions
Recall that Écalle’s resurgent series have the special property of “reconstructing themselves”. Namely, starting from the germ at the origin, that is, the Borel transform of the divergent formal power series, and analytically continuing it to the whole complex plane, it is possible to recover the germs at the other singularities. We refer to Section 2.1 for an introduction to the theory of resurgence. In the example of the spectral trace of local ${\mathbb{P}}^{2}$, the resurgent structure appears trivial at first glance, as the germs at the other singularities are all constants. Yet, as discovered in [18] and fully outlined in [19], when regarded as a sequence, these constants possess rich analytic number-theoretic properties that make contact with the notion of quantum modularity. Furthermore, when considering both weak and strong coupling limits, a unique strong-weak resurgent symmetry holds. As they can be generalized to a whole class of asymptotic series, the results of [19] encourage us to adopt a new perspective and motivate the following discussion.
We will consider a particular type of resurgent series and present a new paradigm of resurgence built upon the Stokes constants, their generating functions, and their Dirichlet series. We refer to this new paradigm as modular resurgence for reasons that will become clear in Section 3.2. Let us start by describing the key ingredients in this construction and the net of relations among them along the same lines of [19, Section 3].
  • Let $\tilde{f}_{\infty}(y)\in{\mathbb{C}}[\![y]\!]$ be a Gevrey-1 formal power series whose resurgent structure consists of a single tower of simple poles at the locations $\zeta_{m}={\cal E}_{0}m$, $m\in{\mathbb{Z}}_{\neq 0}$, for some constant ${\cal E}_{0}\in{\mathbb{C}}$. Let $S_{m}\in{\mathbb{C}}$ be the Stokes constant at the singularity $\zeta_{m}$. By Proposition 3.1, this is equivalent to stating that the perturbative coefficients of $\tilde{f}_{\infty}(y)$ can be written in terms of $\zeta_{m}$ and $S_{m}$ in the form of Eq. (24).
  • Assume that the complex numbers $S_{m}$, $m\in{\mathbb{Z}}_{\neq 0}$, are the coefficients of two $L$-series $L_{0,+}(s)$ and $L_{0,-}(s)$, where $s\in{\mathbb{C}}$ and $\Re(s)>\alpha\geq 0$, defined as in Eq. (29). Following Proposition 3.2, we can build their generating function $f_{0}(y)$, $y\in{\mathbb{C}}\setminus{\mathbb{R}}$, defined as in Eq. (30), by taking the inverse Mellin transform of $L_{0,\pm}(s)$ as in Eq. (32).
  • Let $\tilde{f}_{0,\pm}(y)\in{\mathbb{C}}[\![y]\!]$ be the asymptotic series of $f_{0}(y)$ as $y\to 0$ with $\Im(y)>0$ and $\Im(y)<0$, respectively. Following Proposition 3.3, the perturbative coefficients of $\tilde{f}_{0,\pm}(y)$ are explicitly given by the values of the $L$-function $L_{0,\pm}(s)$ at negative integers as in Eq. (36). Note that $L_{0,\pm}(-n)$ for $n\in{\mathbb{Z}}_{>0}$ are computed through the functional equations that give the analytic continuation of the corresponding $L$-series $L_{0,\pm}(s)$. Sometimes, the analytic continuation of an $L$-function is described by another $L$-function with different coefficients. Let us assume this is the case and denote the new coefficients by $R_{m}\in{\mathbb{C}}$, $m\in{\mathbb{Z}}_{\neq 0}$.
  • Crucially, the resurgence structures of $\tilde{f}_{0,\pm}(y)$ are completely determined by the functional equations satisfied by $L_{0,\pm}(s)$, respectively. We assume that the perturbative coefficients of $\tilde{f}_{0}(y):=\tilde{f}_{0,+}(y)-\tilde{f}_{0,-}(y)$ can be written in terms of $\eta_{m}$ and $R_{m}$ in the form of Eq. (24), where $\eta_{m}={\cal E}_{\infty}m$, $m\in{\mathbb{Z}}_{\neq 0}$, for some constant ${\cal E}_{\infty}\in{\mathbb{C}}$. Then, by Proposition 3.1, this is equivalent to imposing that $\tilde{f}_{0}(y)$ is a simple resurgent series whose Borel transform has a single tower of simple poles at $\eta_{m}$ with Stokes constants $R_{m}$.
  • Let $L_{\infty,\pm}(s)$, where $s\in{\mathbb{C}}$ and $\Re(s)>\alpha\geq 0$, be the two $L$-series, defined as in Eq. (29), with coefficients $R_{m}$, $m\in{\mathbb{Z}}_{\neq 0}$. Following Proposition 3.2, we can build their generating function $f_{\infty}(y)$, $y\in{\mathbb{C}}\setminus{\mathbb{R}}$, defined as in Eq. (30), by taking the inverse Mellin transform of $L_{\infty,\pm}(s)$ as in Eq. (32).
  • Finally, let $\tilde{f}_{\infty,\pm}(y)\in{\mathbb{C}}[\![y]\!]$ be the asymptotic series of $f_{\infty}(y)$ as $y\to 0$ with $\Im(y)>0$ and $\Im(y)<0$, respectively. Again following Proposition 3.3, the perturbative coefficients of $\tilde{f}_{\infty,\pm}(y)$ are explicitly given by the values of the $L$-function $L_{\infty,\pm}(s)$ at negative integers as in Eq. (36). Once more, $L_{\infty,\pm}(-n)$ for $n\in{\mathbb{Z}}_{>0}$ are computed through the functional equations that give the analytic continuation of the corresponding $L$-series $L_{\infty,\pm}(s)$. By construction, these are determined by $L_{0,\pm}(s)$. We recover $\tilde{f}_{\infty}(y)$ as the combination $\tilde{f}_{\infty}(y)=\tilde{f}_{\infty,+}(y)-\tilde{f}_{\infty,-}(y)$.
Notice that the above relations among the dual resurgent structures of the pair of asymptotic series $\tilde{f}_{0}$, $\tilde{f}_{\infty}$ assemble into a unique global construction. We summarize the previous circular argument in the commutative diagram below.
[…]
Notably, starting from any fixed vertex, all other vertices of this symmetric diagram are spanned by following its directed arrows – that is, the information contained in each vertex reconstructs the whole diagram and is therefore equivalent to the information content of every other vertex. Moreover, a special place is occupied by the two-headed diagonal arrow representing the functional equation connecting the two pairs of $L$-functions $L_{0,\pm}(s)$ and $L_{\infty,\pm}(s)$.
Let us state and prove the three propositions on which the modular resurgence paradigm is built. The following result is a generalization of the exact large-order relations of [18].
Proposition 3.1.
Let $\tilde{f}(y)=\sum_{n=1}^{\infty}c_{n}y^{n}\in{\mathbb{C}}[\![y]\!]$ be a Gevrey-1 asymptotic series with coefficients given explicitly by
$c_{n}=\frac{\Gamma(n)}{2\pi{\rm i}}\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{\zeta_{m}^{n}}\,,\quad n\in{\mathbb{Z}}_{>0}\,,$ (24)
where $A_{m}$, $\zeta_{m}$, $m\in{\mathbb{Z}}_{\neq 0}$, are complex numbers. Its Borel transform is a simple resurgent function with simple poles at the locations $\zeta_{m}$ and Stokes constants $A_{m}$.
Proof.
By definition, the Borel transform of $\tilde{f}$ is the function
${\cal B}[\tilde{f}](\zeta)=\sum_{n=1}^{\infty}\frac{c_{n}}{\Gamma(n)}\zeta^{n-1}=\frac{1}{2\pi{\rm i}}\sum_{n=1}^{\infty}\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{\zeta_{m}^{n}}\zeta^{n-1}\in{\mathbb{C}}\{\zeta\}\,.$ (25)
Since $\tilde{f}$ is Gevrey-1, the coefficients $c_{n}$ have the factorially divergent growth
$|c_{n}|\leq n!|{\cal E}|^{-n}\quad n\gg 1\,,\quad{\cal E}\in{\mathbb{C}}\,.$ (26)
Hence, we can change the order of summation in Eq. (25), giving
${\cal B}[\tilde{f}](\zeta)=\frac{1}{2\pi{\rm i}}\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{\zeta_{m}}\sum_{n=1}^{\infty}\left(\frac{\zeta}{\zeta_{m}}\right)^{n-1}=-\frac{1}{2\pi{\rm i}}\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{\zeta-\zeta_{m}}\,,$ (27)
which has simple poles at $\zeta=\zeta_{m}$ with corresponding Stokes constants $A_{m}$. ∎
Note that Eq. (24) can be equivalently written as
$c_{n}=\frac{\Gamma(n)}{(2\pi{\rm i}){\cal E}^{n}}\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{m^{n}}\,,\quad n\in{\mathbb{Z}}_{>0}\,,$ (28)
when $\zeta_{m}={\cal E}m$, $m\in{\mathbb{Z}}_{\neq 0}$, for some constant ${\cal E}\in{\mathbb{C}}$. Up to the prefactors, the right-hand-side (RHS) in Eq. (28) is then the Dirichlet series with coefficients given by the Stokes constants $A_{m}$ evaluated at the positive integer $n$. This does not a priori define an $L$-series. However, when it does, we can say something more.
Proposition 3.2.
Let $\{A_{m}\}$, $m\in{\mathbb{Z}}_{\neq 0}$, be a sequence of complex numbers defining the $L$-series
$L_{+}(s)=\sum_{m>0}\frac{A_{m}}{m^{s}}\,,\quad L_{-}(s)=-\sum_{m>0}\frac{A_{-m}}{m^{s}}$ (29)
that are convergent in some right half-plane $\{\Re(s)>\alpha\}\subset{\mathbb{C}}$, $\alpha\geq 0$. Let $f\colon{\mathbb{C}}\setminus{\mathbb{R}}\to{\mathbb{C}}$ be the generating function
$f(y):=\begin{cases}\displaystyle\sum_{m>0}A_{m}{\rm e}^{2\pi{\rm i}my}&\text{ if }\Im(y)>0\\ &\\ -\displaystyle\sum_{m<0}A_{m}{\rm e}^{2\pi{\rm i}my}&\text{ if }\Im(y)<0\end{cases}\,.$ (30)
There is a canonical correspondence between the pair of $L$-series $L_{+}(s),L_{-}(s)$ and the generating function $f(y)$.
Proof.
The statement follows from the properties of the Mellin transform. Indeed, a simple computation shows that
[…]
that is, the $L$-series $L_{\pm}(s)$ are given by the Mellin transform of the generating function $f(y)$. Conversely, by the Mellin inversion theorem, $f(y)$ can be computed via the inverse Mellin transform of $L_{\pm}(s)$. Indeed, for $C>\alpha$, we have that
[…]
which can be analytically continued to all arguments $y\in{\mathbb{C}}\setminus{\mathbb{R}}$. ∎
Let us stress that the generating function $f(y)$ in Eq. (30) has the form of a $q$-series for $q={\rm e}^{2\pi{\rm i}y}$ when restricted to either the upper or the lower halves of the complex plane. Besides, we could equivalently consider another pair of $L$-series, such as
$L(s)=\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{|m|^{s}}\,,\quad L^{\prime}(s)=\sum_{m\in{\mathbb{Z}}_{\neq 0}}\mathrm{sgn}(m)\frac{A_{m}}{|m|^{s}}\,,$ (33)
which are again in a one-to-one correspondence with the generating function $f(y)$ in Eq. (30) as a consequence of Proposition 3.2. Indeed, $L(s)$ and $L^{\prime}(s)$ can be written as the linear combinations
$L(s)$ $=\sum_{m>0}\frac{A_{m}}{m^{s}}+\sum_{m<0}\frac{A_{m}}{(-m)^{s}}=L_{+}(s)-L_{-}(s)\,,$ (34)
$L^{\prime}(s)$ $=\sum_{m>0}\frac{A_{m}}{m^{s}}-\sum_{m<0}\frac{A_{m}}{(-m)^{s}}=L_{+}(s)+L_{-}(s)\,.$
We can then consider the meromorphic extension of these $L$-series – or the $L$-series in Eq. (29), to the whole complex $s$-plane, that is, the corresponding $L$-functions. The following result states that knowing in closed form the perturbative coefficients of the asymptotic series obtained by expanding the generating function $f(y)$ at zero for $\Im(y)>0$ is equivalent to computing the analytic continuation of the $L$-series $L_{+}(s)$. An analogous statement holds for $\Im(y)<0$.
Proposition 3.3.
Let $\{A_{m}\}$, $m\in{\mathbb{Z}}_{>0}$, be a sequence of complex numbers defining an $L$-series
$L(s)=\sum_{m=1}^{\infty}\frac{A_{m}}{m^{s}}$ (35)
that is convergent in some right half-plane $\{\Re(s)>\alpha\}\subset{\mathbb{C}}$, $\alpha\geq 0$, and can be analytically continued to $\{\Re(s)<0\}$ through a functional equation. The following statements apply.
  • 1.
    The generating function $f(y):=\sum_{m=1}^{\infty}A_{m}{\rm e}^{2\pi{\rm i}my}$ converges for $y\in{\mathbb{H}}$.
  • 2.
    Let $\tilde{f}(y)=\sum_{n=0}^{\infty}c_{n}y^{n}\in{\mathbb{C}}[\![y]\!]$ be the asymptotic series of $f(y)$ as $y\to 0$ with $\Im(y)>0$. The perturbative coefficients $c_{n}$, $n\in{\mathbb{Z}}_{\geq 0}$, are given by
    $c_{n}=L(-n)\frac{(2\pi{\rm i})^{n}}{n!}\,,$ (36)
    where $L(-n)$ is defined by the functional equation.
Proof.
As we have seen in Proposition 3.2, the generating function $f(y)$ is the inverse Mellin transform of $(2\pi)^{-s}\Gamma(s)L(s)$. Indeed, let $t>0$ and fix $C>\alpha$. Then,
$f({\rm i}t)$ $=\frac{1}{2\pi{\rm i}}\int_{C-{\rm i}\infty}^{C+{\rm i}\infty}(2\pi t)^{-s}\Gamma(s)\sum_{m=1}^{\infty}\frac{A_{m}}{m^{s}}\,ds$ (37)
[…]
which can be analytically continued to a function $f(y)$ for $y\in{\mathbb{H}}$. Taking the power series expansion of the exponential ${\rm e}^{2\pi{\rm i}my}$ at $y=0$, the generating function gives
$f(y)=\sum_{m=1}^{\infty}A_{m}{\rm e}^{2\pi{\rm i}my}\sim$ $\sum_{m=1}^{\infty}A_{m}\sum_{n=0}^{\infty}\frac{(2\pi{\rm i})^{n}}{n!}m^{n}y^{n}$ (38)
``="````="` ` = " $\sum_{n=0}^{\infty}\sum_{m=1}^{\infty}\frac{A_{m}}{m^{-n}}\frac{(2\pi{\rm i})^{n}}{n!}y^{n}$
$=$ $\sum_{n=0}^{\infty}L(-n)\frac{(2\pi{\rm i})^{n}}{n!}y^{n}\,,$
where we have changed the order of summation. [^6] The change of summation order is a formal operation here as the coefficients $c_{n}$ might be factorially divergent. Finally, we make sense of $L(-n)$ through the analytic continuation of $L(s)$. ∎
Remark 3.1.
Depending on the explicit form of the functional equation satisfied by the $L$-function $L(s)$ in Eq. (35), it might happen that $L(-n)$ is not well-defined for some $n\in{\mathbb{Z}}_{\geq 0}$. In this case, the computation of the corresponding perturbative coefficients $c_{n}$ through the formula in Eq. (36) requires a separate ad-hoc analysis. A complete first-principles understanding of the domain of validity of Eq. (36) necessitates a specific investigation of the role played by the parameters entering the functional equation, its central point, and the motivic weight of the $L$-function. We leave this endeavor for the future.
Let us highlight how Proposition 3.3 implies that the asymptotic series of the generating function of the Stokes constants is governed by the functional equation satisfied by the corresponding $L$-function. This is the core of the modular resurgence paradigm and explains why, starting from a given $L$-function, one might find a new divergent formal power series with a new set of Stokes constants whose resurgent structure is, however, determined by the original $L$-function. In some examples, the analytic continuation of an $L$-function is given by the $L$-function itself, in which case modular resurgence simplifies.
Remark 3.2.
As we have pointed out, the modular resurgence paradigm is governed by the functional equation for the $L$-functions defined by the Stokes constants. It might happen that the functional equations for the $L$-functions $L_{\pm}(s)$ in Eq. (29) is written in terms of the same $L$-functions – as is the case for the Riemann zeta function $\zeta(s)$, for example. Equivalently, it might happen that the generating function $f(y)$ of a set of Stokes constants $\{A_{m}\}$, $m\in{\mathbb{Z}}_{\neq 0}$, which are obtained by the resurgence of the Gevrey-1 formal power series $\tilde{f}_{+}(y)$ and $\tilde{f}_{-}(y)$, has asymptotic series in the limit $y\to 0$ with $\Im(y)>0$ (respectively, $\Im(y)<0$) that are equal to $\tilde{f}_{\pm}(y)$. In this case, the commutative diagram in Eq. (23) reduces to the one below.
[…]
This simplified version of the modular resurgence construction is observed, among others, in the examples of the series $\sigma$ and $\sigma^{*}$, whose $L$-function was first studied by Cohen in [30]. [^7] The study of the resurgent structure of $\sigma$ and $\sigma^{*}$ was presented in [23, from min. 21.55]. From a number-theoretic point of view, these examples were studied by Zagier in relation to quantum modular forms [6]. More generally, a similar behavior can be observed when $f(y)$ is the function associated with a Maass cusp form of the kind studied in [31]. See Proposition 4.1.
From the conventional point of view in resurgence, looking at the generating series of the Stokes constants is unusual, as one would not generally expect them to define a function (except in those examples where there are finitely many Stokes constants, whose generating series is indeed a polynomial). However, we showed that when the resurgent structure is simple enough, and the Stokes constants have special analytic and number-theoretic properties, studying their generating functions allows us to construct a new formal power series whose resurgent structure is deeply related to one of the original formal power series. In particular, when the Stokes constants are coefficients of an $L$-function, the modular resurgence scheme applies and is governed by the functional equation. Therefore, we advocate a new perspective on resurgence that aims to investigate the properties of the Stokes constants themselves.
3.2 Modular resurgent structures
Our previous discussion highlights the leading part played by the $L$-functions and their functional equation in describing new features of a particular type of resurgent series. In this section, we address the relationship between quantum modularity and resurgent structures, which will clarify our choice of name for the paradigm of modular resurgence. We start by proposing the following definitions of modular resurgent structures and modular resurgent series.
Definition 3.1.
A Gevrey-1 asymptotic series $\tilde{f}(y)\in{\mathbb{C}}\llbracket y\rrbracket$ has a modular resurgent structure if the following conditions hold.
  1. 1.
    The Borel transform ${\cal B}[\tilde{f}](\zeta)\in{\mathbb{C}}\{\zeta\}$ has a tower of singularities at the locations $\zeta_{m}={\cal E}m$, $m\in\mathbb{Z}_{\neq 0}$, for some constant ${\cal E}\in{\mathbb{C}}$, in the complex $\zeta$-plane.
  2. 2.
    For every $m\in{\mathbb{Z}}_{\neq 0}$, the resurgent series at the singularity $\zeta_{m}$ is the constant function $A_{m}\in{\mathbb{C}}$, i.e., the Stokes constant.
  3. 3.
    The Stokes constants $A_{m}$, $m\in{\mathbb{Z}}_{\neq 0}$, are the coefficients of two $L$-series
    $L(s)=\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{|m|^{s}}\,,\quad L^{\prime}(s)=\sum_{m\in{\mathbb{Z}}_{\neq 0}}\mathrm{sgn}(m)\frac{A_{m}}{|m|^{s}}\,,$ (40)
    that are convergent in some right half-plane $\{\Re(s)>\alpha\}\subset{\mathbb{C}}$, $\alpha\geq 0$, and can be analytically continued to $\Re(s)<0$.
Definition 3.2.
A Gevrey-1 asymptotic series with a modular resurgent structure is a modular resurgent series. The Borel transform of a modular resurgent series is a modular resurgent function.
It follows from Proposition 3.2 that a modular resurgent series is equivalently characterized by the $L$-functions $L_{\pm}(s)$ in Eq. (29) or the generating function $f(y)$ in Eq. (30). Moreover, the discontinuities across the Stokes rays identified by the tower of singularities are given by
$\mathrm{disc}_{\theta}\,\tilde{f}_{+}(y)$ $=\sum_{m=1}^{\infty}A_{m}\,{\rm e}^{{\cal E}my}=f\bigg{(}\frac{{\cal E}y}{2\pi{\rm i}}\bigg{)}\,,\quad y\in{\mathbb{H}}\cap\{\Re(e^{-{\rm i}\theta}y)>0\}\,,$ (41a)
$\mathrm{disc}_{2\pi-\theta}\,\tilde{f}_{-}(y)$ $=\sum_{m=1}^{\infty}A_{-m}\,{\rm e}^{-{\cal E}my}=-f\bigg{(}\frac{{\cal E}y}{2\pi{\rm i}}\bigg{)}\,,\quad y\in{\mathbb{H}}_{-}\cap\{\Re(e^{-{\rm i}\theta}y)>0\}\,,$ (41b)
where $0\leq\theta<\pi$ is the argument of the singularities in the upper half of the Borel plane and $\tilde{f}_{\pm}(y)$ are the asymptotic series of $f(y)$ as $y\to 0$ with $\Im(y)>0$ and $\Im(y)<0$, respectively.
Let us now state our main conjectures on modular resurgence. A $q$-series $f\colon{\mathbb{H}}\to{\mathbb{C}}$ whose asymptotic series at zero has a modular resurgent structure is expected to agree with the median resummation of its asymptotic expansion and to be a holomorphic quantum modular form.
Conjecture 1.
Let $f\colon{\mathbb{H}}\to{\mathbb{C}}$ be a $q$-series where $q={\rm e}^{2\pi{\rm i}y}$. If its asymptotic expansion $\tilde{f}(y)$ as $y\to 0$ with $\Im(y)>0$ has a modular resurgent structure, then the median resummation of $\tilde{f}(y)$ reconstructs the original function $f(y)$, that is,
$\mathcal{S}_{\theta}^{\mathbf{med}}\tilde{f}(y)=f(y)\,,\quad y\in{\mathbb{H}}\cap\{\Re(e^{-{\rm i}\theta}y)>0\}\,,$ (42)
where $\theta$ is the argument of the singularities in the Borel plane.
Conjecture 2.
Let $f\colon{\mathbb{H}}\to{\mathbb{C}}$ be a $q$-series where $q={\rm e}^{2\pi{\rm i}y}$. If its asymptotic expansion $\tilde{f}(y)$ as $y\to 0$ with $\Im(y)>0$ has a modular resurgent structure, then the function $f(y)$ is a holomorphic quantum modular form for a subgroup $\Gamma\subseteq\mathsf{SL}_{2}(\mathbb{Z})$.
We can relax the assumptions in the conjectures above as for Zagier’s strange identity [20]. Namely, let $f^{\prime}\colon{\mathbb{Q}}\to{\mathbb{C}}$ be such that there exists a $q$-series $f\colon{\mathbb{H}}\to{\mathbb{C}}$ whose radial limit at the non-zero rationals agrees with $f^{\prime}$ to all orders. If the asymptotic expansion $\tilde{f}(y)$ as $y\to 0$ with $\Im(y)>0$ has a modular resurgent structure, then we expect that the median resummation of $\tilde{f}(y)$ restricted to the non-zero rationals reconstructs the original function $f^{\prime}(y)$, that is,
$\mathcal{S}_{\theta}^{\mathbf{med}}\tilde{f}(y)=f^{\prime}(y)\,,\quad y\in{\mathbb{Q}}\cap\{\Re(e^{-{\rm i}\theta}y)>0\}\,,$ (43)
where $\theta$ is again the argument of the singularities in the Borel plane. At the same time, we expect that the function $f^{\prime}(y)$ is a quantum modular form for a subgroup $\Gamma\subseteq\mathsf{SL}_{2}(\mathbb{Z})$.
Remark 3.3.
Note that the generating function of the Stokes constants defined in Eq. (30) is originally supported on ${\mathbb{C}}\setminus{\mathbb{R}}$. Indeed, all statements above can be reformulated for a $q$-series $f\colon{\mathbb{H}}_{-}\to{\mathbb{C}}$, where ${\mathbb{H}}_{-}$ denotes the lower half of the complex plane.
Furthermore, we can consider functions that are defined only over the rationals. This was, in fact, the setting where quantum modular forms were originally introduced [6], as briefly discussed in Section 2.2. Conjectures 1 and 2 then become the following.
Conjecture 3.
Let $f\colon{\mathbb{Q}}\to{\mathbb{C}}$ be a $q$-series where $q={\rm e}^{2\pi{\rm i}y}$. If its asymptotic expansion $\tilde{f}(y)$ as $y\to 0$ extends to $y\in{\mathbb{C}}$ and has a modular resurgent structure, then the median resummation of $\tilde{f}(y)$ reconstructs the original function $f(y)$, that is,
$\mathcal{S}_{\theta}^{\mathbf{med}}\tilde{f}(y)=f(y)\,,\quad y\in{\mathbb{Q}}\cap\{\Re(e^{-{\rm i}\theta}y)>0\}\,,$ (44)
where $\theta$ is the argument of the singularities in the Borel plane.
Conjecture 4.
Let $f\colon{\mathbb{Q}}\to{\mathbb{C}}$ be a $q$-series where $q={\rm e}^{2\pi{\rm i}y}$. If its asymptotic expansion $\tilde{f}(y)$ as $y\to 0$ extends to $y\in{\mathbb{C}}$ and has a modular resurgent structure, then the function $f(y)$ is a quantum modular form for a subgroup $\Gamma\subseteq\mathsf{SL}_{2}(\mathbb{Z})$.
Note that, to perform a resurgent analysis of $\tilde{f}(y)$ in Conjectures 3 and 4, it is necessary to complexify the variable $y$. Namely, $\tilde{f}(y)$ must be analytically continued to $y\in{\mathbb{C}}$. In Conjecture 3, the median resummation $\mathcal{S}_{\theta}^{\mathbf{med}}\tilde{f}(y)$ is then restricted to the original domain of the function $f(y)$.
Remark 3.4.
At this stage, other assumptions may be needed to fully characterize the $L$-function associated with the generating function of the Stokes constants and enable us to prove the conjectures above within the paradigm of modular resurgence. In particular, we expect the functional equation for the $L$-function to be a key ingredient. A complete characterization of the $L$-function would require assumptions on the gamma factors appearing in its completion. We plan to probe these ideas in the future.
We conclude by listing the currently available evidence for the conjectures above.
  • In support of Conjectures 1 and 2, we have the generating functions of the weak and strong coupling Stokes constants for $\log\mathrm{Tr}(\mathsf{\rho}_{{\mathbb{P}}^{2}})$ [19] (see also Section 4.1) and a collection of examples built from the theory of Maass cusp forms and their periods (see Section 4.2).
  • In support of Conjecture 3, we have the Kontsevich–Zagier series [^8] In the example of the Kontsevich–Zagier series, the resurgent structure has a tower of fractional power singularities, and not simple poles, contrary to what occurs in the explicit examples discussed in this paper. [1] and the WRT invariants for a family of Seifert fibered homology spheres [2, Theorem 3]. Besides, according to [1, Conjecture 1.4], more examples might come from the quantum invariants of knots and 3-manifolds.
  • In support of Conjecture 4, we cite again the Kontsevich–Zagier series, whose quantum modularity was proven in [20], the LMO ($\hat{Z}$) invariants for a family of Seifert fibered homology spheres [13, 14], and the $q$-series $\sigma$ and $\sigma^{*}$, which are quantum modular forms according to [6, Example 1].
Note that not all (holomorphic) quantum modular forms arise from modular resurgent series. For instance, the function $\tilde{\Delta}$ in Eq. (20) is not an example of Conjecture 2 because its asymptotic series is a polynomial, hence its resurgent structure is trivial.
Moreover, among the examples providing evidence of the Conjectures 1 and 2, we did not mention the quantum invariants of hyperbolic knots, except for the resurgent conjecture in [1, Conjecture 1.4]. The Kashaev invariants of hyperbolic knots are conjecturally vector-valued quantum modular forms [7]. However, their resurgent structure does not display a single tower of singularities but the so-called peacock pattern [32, 33]. Peacock patterns are also conjectured to appear in the resurgence of the fermionic spectral traces of toric CY threefolds according to [34, 18]. It would be interesting to investigate examples exhibiting multiple towers of singularities in the Borel plane and possibly extend our definition of modular resurgent structures accordingly. We plan to explore this direction in the future.
4 Examples of modular resurgence
In this section, we describe two complete examples of the new paradigm of modular resurgence. In particular, we show how the global exact resurgent symmetry of Section 3.1 reproduces the strong-weak symmetry discovered in [19, Section 3] once specialized to the spectral trace of local ${\mathbb{P}}^{2}$. Then, we prove our conjectural statements in Section 3.2 for a large class of modular resurgent series originating from the theory of Maass cusp forms. We use the notation ${\mathbb{C}}^{\prime}={\mathbb{C}}\setminus{\mathbb{R}}_{\leq 0}$.
4.1 From topological string theory
Let us compare the paradigm of modular resurgence with the strong-weak resurgent symmetry of the spectral trace of local ${\mathbb{P}}^{2}$, showing that the construction presented here and the one of [19, Section 3] are coincident and complementary.
We start by recalling the necessary background from [18]. Following the prescription of [35, 36], quantizing the mirror curve to local ${\mathbb{P}}^{2}$ gives rise to a quantum operator acting on $L^{2}({\mathbb{R}})$ whose inverse $\rho_{{\mathbb{P}}^{2}}$ is positive-definite and of trace class. Its first fermionic spectral trace [^9] The fermionic spectral traces $Z_{X}(\bm{N},\hbar)$ provide a non-perturbative definition of the topological string on a toric CY threefold $X$ according to the conjectural Topological String/Spectral Theory (TS/ST) correspondence. The Planck constant satisfies a strong-weak coupling duality $\hbar\propto g_{s}^{-1}$, where $g_{s}$ is the string coupling constant. $Z_{{\mathbb{P}}^{2}}(1,\hbar)=\mathrm{Tr}(\mathsf{\rho}_{{\mathbb{P}}^{2}})$ is a well-defined analytic function of the quantum deformation parameter $\hbar\in{\mathbb{C}}^{\prime}$, which is known in closed form for $\Im(\hbar)>0$ [37]. The all-orders perturbative expansions of the logarithm of the spectral trace in the limits of $\hbar\rightarrow 0$ and $\hbar\rightarrow\infty$ have been computed in [18]. They give rise to the Gevrey-1 asymptotic series
ϕ()italic-$\phi(\hbar)$ $=\sum_{n=1}^{\infty}a_{2n}\hbar^{2n}\in{\mathbb{Q}}[\![\hbar]\!]\,,$ (45a)
$\psi(\tau)$ $=\sum_{n=1}^{\infty}b_{2n}\tau^{2n-1}\in{\mathbb{Q}}[\pi,\sqrt{3}][\![\tau]\!]\,,$ (45b)
where $\tau=-\tfrac{2\pi}{3\hbar}$, respectively. It is proven in [18] that their Borel transforms possess a single tower of simple poles at the locations
$\zeta_{m}={\cal A}_{0}{\rm i}m\,,\quad\eta_{m}={\cal A}_{\infty}{\rm i}m\,,\quad m\in{\mathbb{Z}}_{\neq 0}\,,$ (46)
where ${\cal A}_{0}=\frac{4\pi^{2}}{3}$ and ${\cal A}_{\infty}=\frac{2\pi}{3}$, respectively. The corresponding Stokes constants
$S_{m}=S_{-m}\in\sqrt{3}{\rm i}\,{\mathbb{Q}}_{>0}\,,\quad R_{m}=-R_{-m}\in{\mathbb{Q}}_{\neq 0}\,,\quad m\in{\mathbb{Z}}_{\neq 0}\,,$ (47)
are known explicitly as divisor sum functions and define the weak and strong coupling $L$-series
$L_{0}(s)=\sum_{m=1}^{\infty}\frac{S_{m}}{m^{s}}\,,\quad L_{\infty}(s)=\sum_{m=1}^{\infty}\frac{R_{m}}{m^{s}}\,,$ (48)
which converge in the right half-plane $\Re(s)>1$ and can be analytically continued to meromorphic functions on the whole complex $s$-plane [18, Section 4.4]. Therefore, ϕ()italic-$\phi(\hbar)$ and $\psi(\tau)$ are modular resurgent series according to Definition 3.1.
In the companion paper [19], we complete the $L$-functions in Eq. (48) into
$\Lambda_{0}(s)$ $=\frac{3^{\frac{s}{2}-1}}{{\rm i}\pi^{s+1}}\Gamma\left(\frac{s}{2}\right)\Gamma\left(\frac{s}{2}+1\right)L_{0}(s)\,,$ (49a)
$\Lambda_{\infty}(s)$ $=\frac{3^{\frac{s}{2}-1}}{\pi^{s+1}}\Gamma\left(\frac{s+1}{2}\right)^{2}L_{\infty}(s)\,,$ (49b)
and show that $\Lambda_{0}(s)$, $\Lambda_{\infty}(s)$ are analytically continued to the whole complex $s$-plane through each other as a consequence of the factorization formulae of [18]. In particular, they satisfy the functional equation
$\Lambda_{0}(s)=\Lambda_{\infty}(-s)\,.$ (50)
Simultaneously, we take the perspective of the generating functions $f_{0}(y)$, $f_{\infty}(y)$, $y\in{\mathbb{C}}\setminus{\mathbb{R}}$, of the weak and strong coupling Stokes constants, which satisfy the parity properties
$f_{0}(-y)=-f_{0}(y)\,,\quad f_{\infty}(-y)=f_{\infty}(y)\,.$ (51)
As a consequence of the discontinuity formulae of [18], we obtain that the all-orders perturbative expansions in the limit $y\rightarrow 0$ with $\Im(y)>0$ of the generating functions as
$\tilde{f}_{0}(y)$ $=-\frac{\pi{\rm i}}{2}-\frac{3\mathcal{V}}{2\pi{\rm i}y}-2\psi(y)\,,$ (52a)
$\tilde{f}_{\infty}(y)$ $=-3\log\frac{\Gamma(2/3)}{\Gamma(1/3)}-\log(-6\pi{\rm i}y)+2\phi(2\pi y)\,,$ (52b)
where $\mathcal{V}=2\Im\left(\mathrm{Li}_{2}({\rm e}^{2\pi{\rm i}/3})\right)$.
Let us now neglect the terms that do not enter the perturbative series in the RHS of Eqs. (52a) and (52b). We show that they satisfy the newly introduced paradigm of modular resurgence. This offers a complementary understanding of the strong-weak resurgent symmetry discussed in [19, Section 3], particularly highlighting the central role played by the functional equation in Eq. (50). For completeness, we repeat the main steps in the modular resurgence construction for this key example.
  • The resurgent structure of the asymptotic series $\tilde{f}_{\infty}(y)$ is dictated by the formal power series $2\phi(2\pi y)$ according to the formula in Eq. (52b). Thus, the resurgent structure is modular and can be written in the coordinate conjugate to $y$ as $\{\eta_{m},\,S_{m}\}$, $m\in{\mathbb{Z}}_{\neq 0}$, where the Stokes constants satisfy $S_{-m}=S_{m}$.
  • The weak coupling $L$-series $L_{0}(s)$, $s\in{\mathbb{C}}$ with $\Re(s)>1$, has coefficients given by the Stokes constants $S_{m}$ for $m\in{\mathbb{Z}}_{>0}$. Note that $L_{0}(s)=\pm L_{0,\pm}(s)$ in the notation of Eq. (29). Because of the parity of the Stokes constants, we need only consider one $L$-series.
  • Taking the inverse Mellin transform of $L_{0}(s)$ as in Eq. (32), we obtain the generating function $f_{0}(y)$. Let
    $\tilde{f}_{0}(y)=\sum_{n=1}^{\infty}c_{n}^{0}y^{n}\in{\mathbb{C}}\llbracket y\rrbracket$ (53)
    be its asymptotic series as $y\to 0$ with $\Im(y)>0$. By Proposition 3.3, particularly Eq. (36), the perturbative coefficients $c_{n}^{0}$, $n\in{\mathbb{Z}}_{>0}$, are given by
    $c_{n}^{0}=L_{0}(-n)\frac{(2\pi{\rm i})^{n}}{n!}\,,$ (54)
    where we make sense of the RHS via the functional equation for the completed $L$-function $\Lambda_{0}(s)$. In particular, substituting Eqs. (49a) and (49b) into Eq. (50), we find that
    $L_{0}(-s)=-\frac{2{\rm i}3^{s}}{\pi^{2s}s}\frac{\Gamma\left(\frac{s+1}{2}\right)^{2}}{\Gamma\left(-\frac{s}{2}\right)^{2}}L_{\infty}(s)\,.$ (55)
    Notice that $L_{0}(-s)$ tends to zero for $s=n\in{\mathbb{Z}}_{>0}$ even as a consequence of the divergence of the gamma function in the denominator. Thus, we set $c_{2n}^{0}=0$, $n\in{\mathbb{Z}}_{>0}$. Then, taking $s=n\in{\mathbb{Z}}_{>0}$ odd, applying the well-known properties of the gamma function, and using the series representation of $L_{\infty}(s)$ in Eq. (48), Eq. (55) becomes
    $L_{0}(-n)$ $=-\frac{2{\rm i}}{\pi}n!(n-1)!(-2\pi{\rm i})^{-n}\sum_{m=1}^{\infty}\frac{R_{m}}{\eta_{m}^{n}}\,.$ (56)
    Applying Eq. (54), the non-trivial perturbative coefficients $c_{2n+1}^{0}$, $n\in{\mathbb{Z}}_{\geq 0}$, are
    $c_{2n+1}^{0}=-2\frac{\Gamma(2n+1)}{\pi{\rm i}}\sum_{m=1}^{\infty}\frac{R_{m}}{\eta_{m}^{2n+1}}=-2b_{2n+2}\,,$ (57)
    where we have used the exact large-order identity in Eq. (24) for the perturbative coefficients $\{b_{2k}\}$, $k\in{\mathbb{Z}}_{>0}$, of the formal power series $\psi(\tau)$ in Eq. (45b). This is what we expect from the formula in Eq. (52a).
  • The resurgent structure of the asymptotic series $\tilde{f}_{0}(y)$ is dictated by the formal power series $-2\psi(y)$ according to the formula in Eq. (52a). Thus, the resurgent structure is modular and can be written in the coordinate conjugate to $y$ as $\{\eta_{m},R_{m}\}$, $m\in{\mathbb{Z}}_{\neq 0}$, where the Stokes constants satisfy $R_{-m}=-R_{m}$.
  • The strong coupling $L$-series $L_{\infty}(s)$, $s\in{\mathbb{C}}$ with $\Re(s)>1$, has coefficients given by the Stokes constants $R_{m}$ for $m\in{\mathbb{Z}}_{>0}$. Note that $L_{\infty}(s)=L_{\infty,\pm}(s)$ in the notation of Eq. (29). Again, because of the parity of the Stokes constants, we need only consider one $L$-series.
  • Taking the inverse Mellin transform of $L_{\infty}(s)$ as in Eq. (32), we obtain the generating function $f_{\infty}(y)$. Let
    $\tilde{f}_{\infty}(y)=\sum_{n=1}^{\infty}c_{n}^{\infty}y^{n}\in{\mathbb{C}}\llbracket y\rrbracket$ (58)
    be its asymptotic series as $y\to 0$ with $\Im(y)>0$. By Proposition 3.3, particularly Eq. (36), its perturbative coefficients $c_{n}^{\infty}$, $n\in{\mathbb{Z}}_{>0}$, are given by
    $c_{n}^{\infty}=L_{\infty}(-n)\frac{(2\pi{\rm i})^{n}}{n!}\,,$ (59)
    where we make sense of the RHS via the functional equation for the completed $L$-function $\Lambda_{\infty}(s)$ as before. In particular, Eq. (55) gives
    $L_{\infty}(-s)=\frac{3^{s}s}{2{\rm i}\pi^{2s}}\frac{\Gamma\left(\frac{s}{2}\right)^{2}}{\Gamma\left(\frac{1-s}{2}\right)^{2}}L_{0}(s)\,.$ (60)
    Notice that $L_{\infty}(-s)$ tends to zero for $s=n\in{\mathbb{Z}}_{\geq 0}$ odd due to the divergence of the gamma function in the denominator. Hence, we set $c_{2n+1}^{\infty}=0$, $n\in{\mathbb{Z}}_{\geq 0}$. Then, taking $s=n\in{\mathbb{Z}}_{>0}$ even, applying the well-known properties of the gamma function, and using the series representation of $L_{0}(s)$ in Eq. (48), Eq. (60) becomes
    $L_{\infty}(-n)$ $=\frac{2}{{\rm i}\pi}n!(n-1)!(-2\pi)^{n}\sum_{m=1}^{\infty}\frac{S_{m}}{\zeta_{m}^{n}}\,.$ (61)
    Applying Eq. (59), the non-trivial perturbative coefficients $c_{2n}^{\infty}$, $n\in{\mathbb{Z}}_{>0}$, are
    $c_{2n}^{\infty}=2(2\pi)^{2n}\frac{\Gamma(2n)}{\pi{\rm i}}\sum_{m=1}^{\infty}\frac{S_{m}}{\zeta_{m}^{2n}}=2(2\pi)^{2n}a_{2n}\,,$ (62)
    where we have used the exact large-order identity in Eq. (24) for the perturbative coefficients $\{a_{2k}\}$, $k\in{\mathbb{Z}}_{>0}$, of the formal power series ϕ()italic-$\phi(\hbar)$ in Eq. (45a). Note again that this is what we expect from the formula in Eq. (52b).
Remark 4.1.
The term of order $y^{0}$ does not enter our definitions of $\tilde{f}_{0}(y)$ and $\tilde{f}_{\infty}(y)$. We observe, however, that direct evaluation of the weak coupling $L$-function $L_{0}(s)$ at $s=0$ using the closed formula in [18, Eq. (4.174a)] gives the perturbative coefficient $c_{0}^{0}=L_{0}(0)=-\pi{\rm i}/2$ according to the formula in Eq. (54), which agrees with the asymptotic expansion in Eq. (52a). At the same time, the strong coupling $L$-function $L_{\infty}(s)$ in [18, Eq. (4.174b)] is not well-defined at $s=0$ – as it can also be observed from the functional equation in Eq. (60). In particular, following Eq. (59), the perturbative coefficient $c_{0}^{\infty}$ would appear to diverge to infinity. We conclude that the constant term in the asymptotic expansion in Eq. (52b) cannot be consistently derived from the analytic continuation of the $L$-function $L_{\infty}(s)$.
We have shown that the weak and strong coupling resurgent structures of the spectral trace of local ${\mathbb{P}}^{2}$ are a modular pair and satisfy the global symmetry of Section 3.1. Remarkably, as we argue in [19], this construction is responsible for the exchange of the perturbative/nonperturbative contributions to the holomorphic and anti-holomorphic blocks in the well-known factorization of the spectral trace, which can be intuitively traced back to the duality between the worldsheet and WKB contributions to the total grand potential of the topological string on local ${\mathbb{P}}^{2}$. Finally, we stress that both Conjectures 1 and 2 are explicitly proven for the generating function $f_{0}$ [19, Theorems 4.6 and 4.7]. Namely, $f_{0}$ is a holomorphic quantum modular function for the congruence subgroup $\Gamma_{1}(3)\subset\mathsf{SL}_{2}({\mathbb{Z}})$ and is reconstructed by the median resummation of its asymptotic expansion. As for the generating function $f_{\infty}$, Conjecture 2 is proven with respect to the same group $\Gamma_{1}(3)$ [19, Theorem 4.7], while Conjecture 1 is only supported by numerical evidence [19, Conjecture 1].
4.2 From the theory of Maass cusp forms
Given a pair of $L$-series satisfying special functional equations, a large class of examples of modular resurgent series can be derived from this data. In particular, building on the results of [31], we consider $L$-functions related to Maass cusp forms with spectral parameter 1/2121/21 / 2.101010A Maass cusp form with spectral parameter $s\in{\mathbb{C}}$ is a smooth $\mathsf{PSL}_{2}({\mathbb{Z}})$-invariant function $u\colon{\mathbb{H}}\to{\mathbb{C}}$ that is an eigenfunction of the hyperbolic Laplacian $\Delta$ with eigenvalue $s(1-s)$. Namely, $\Delta u=s(1-s)u$. These functions give a basis for $L^{2}$ on ${\mathbb{H}}/\mathsf{PSL}_{2}({\mathbb{Z}})$.
Proposition 4.1.
Let $\{A_{m}\}$, $m\in{\mathbb{Z}}_{\neq 0}$, be a sequence of complex numbers such that the two $L$-series
$L_{\epsilon}(s)=\sum_{m=1}^{\infty}\frac{A_{m}+(-1)^{\epsilon}A_{-m}}{m^{s}}\,,\quad\epsilon=0,1\,,$ (63)
are convergent in some right half-plane $\{\Re(s)>\alpha\}\subset{\mathbb{C}}$, $\alpha\geq 0$. Assume that the corresponding completed $L$-functions
$\Lambda_{\epsilon}(s)=\gamma(s+\epsilon)L_{\epsilon}(s)\,,\quad\text{where}\quad\gamma(s)=\frac{1}{4\pi^{s}}\Gamma\left(\frac{s}{2}\right)^{2}\,,$ (64)
are entire functions of finite order and satisfy the functional equations
$\Lambda_{\epsilon}(1-s)=(-1)^{\epsilon}\Lambda_{\epsilon}(s)\,,\quad\epsilon=0,1\,.$ (65)
Let $f\colon{\mathbb{C}}\setminus{\mathbb{R}}\to{\mathbb{C}}$ be the generating function
$f(y):=\begin{cases}\displaystyle\sum_{m>0}A_{m}{\rm e}^{2\pi{\rm i}my}&\text{ if }\Im(y)>0\\ &\\ -\displaystyle\sum_{m<0}A_{m}{\rm e}^{2\pi{\rm i}my}&\text{ if }\Im(y)<0\end{cases}\,.$ (66)
The asymptotic series $\tilde{f}_{+}(y)$ and $\tilde{f}_{-}(y)$ of $f(y)$ as $y\to 0$ with $\Im(y)>0$ and $\Im(y)<0$, respectively, are modular resurgent series.
Proof.
Let us introduce the two $L$-functions
$L_{+}(s)$ $=\sum_{m=1}^{\infty}\frac{A_{m}}{m^{s}}=\frac{L_{0}(s)+L_{1}(s)}{2}\,,$ (67a)
$L_{-}(s)$ $=-\sum_{m=1}^{\infty}\frac{A_{-m}}{m^{s}}=\frac{L_{1}(s)-L_{0}(s)}{2}\,,$ (67b)
as in Eq. (29). It follows from Proposition 3.3, particularly from Eq. (36), that the asymptotic series of $f(y)$ as $y\to 0$ with $\Im(y)>0$ can be written as
$\tilde{f}_{+}(y)=\sum_{n=0}^{\infty}L_{+}(-n)\frac{(2\pi{\rm i})^{n}}{n!}y^{n}\,.$ (68)
Similarly, its asymptotic series as $y\to 0$ with $\Im(y)<0$ is
$\tilde{f}_{-}(y)=\sum_{n=0}^{\infty}L_{-}(-n)\frac{(2\pi{\rm i})^{n}}{n!}y^{n}\,.$ (69)
We can now apply the functional equation for $\Lambda_{\epsilon}(s)$ in Eq. (65) to make sense of $L_{\pm}(-n)$ for $n\in{\mathbb{Z}}_{\geq 0}$. To do so, let us consider separately the case of $n\in{\mathbb{Z}}_{\geq 0}$ even (respectively, odd) for $L_{1}(-n)$ (respectively, $L_{0}(-n)$) due to the expression for the gamma factor $\gamma(s)$. For $n\in{\mathbb{Z}}_{\geq 0}$ even, we have that
$L_{1}(-n)$ $=-\frac{\gamma(n+2)}{\gamma(1-n)}L_{1}(n+1)$ (70)
$=-\frac{1}{\pi^{2n+1}}\frac{\Gamma\big{(}\frac{n}{2}+1\big{)}^{2}}{\Gamma\big{(}\frac{1-n}{2}\big{)}^{2}}L_{1}(n+1)$
$=-\frac{2^{-2n}}{\pi^{2n+2}}\Gamma(n+1)^{2}L_{1}(n+1)\,,$
where we have used the definition for the gamma factor in Eq. (64) and applied the well-known identities for the gamma function. Similarly, for $n\in{\mathbb{Z}}_{>0}$ odd, we have that
$L_{0}(-n)$ $=\frac{\gamma(n+1)}{\gamma(-n)}L_{0}(n+1)$ (71)
$=\frac{1}{\pi^{2n+1}}\frac{\Gamma\big{(}\frac{n+1}{2}\big{)}^{2}}{\Gamma\big{(}-\frac{n}{2}\big{)}^{2}}L_{0}(n+1)$
$=\frac{2^{-2n}}{\pi^{2n+2}}\Gamma(n+1)^{2}L_{0}(n+1)\,.$
Since the gamma function has simple poles at the non-positive integers, we set $L_{1}(-n)=0$ for $n\in{\mathbb{Z}}_{>0}$ odd and $L_{0}(-n)=0$ for $n\in{\mathbb{Z}}_{\geq 0}$ even. Therefore, putting together Eqs. (67a), (70), and (71), we find that
$\dfrac{1}{n!}L_{+}(-n)$ =$=\begin{cases}-\dfrac{2^{-2n-1}}{\pi^{2n+2}}n!\,L_{1}(n+1)=-\dfrac{2^{-2n-1}}{\pi^{2n+2}}n!\,\displaystyle\sum_{m=1}^{\infty}\dfrac{A_{m}-A_{-m}}{m^{n+1}}&\text{ if }n\in{\mathbb{Z}}_{\geq 0}\text{ even}\\ &\\ \dfrac{2^{-2n-1}}{\pi^{2n+2}}n!\,L_{0}(n+1)=\dfrac{2^{-2n-1}}{\pi^{2n+2}}n!\,\displaystyle\sum_{m=1}^{\infty}\dfrac{A_{m}+A_{-m}}{m^{n+1}}&\text{ if }n\in{\mathbb{Z}}_{>0}\text{ odd}\end{cases}$ (72)
$=\frac{2^{-2n-1}}{\pi^{2n+2}}n!\left(\sum_{m=1}^{\infty}\frac{A_{m}}{(-m)^{n+1}}+\sum_{m=1}^{\infty}\frac{A_{-m}}{m^{n+1}}\right)$
$=\frac{2^{-2n-1}}{\pi^{2n+2}}n!\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{(-m)^{n+1}}\,.$
Analogously, putting together Eqs. (67b), (70), and (71), we find that
$\dfrac{1}{n!}L_{-}(-n)$ =$=\begin{cases}-\dfrac{2^{-2n-1}}{\pi^{2n+2}}n!\,L_{1}(n+1)=-\dfrac{2^{-2n-1}}{\pi^{2n+2}}n!\,\displaystyle\sum_{m=1}^{\infty}\dfrac{A_{m}-A_{-m}}{m^{n+1}}&\text{ if }n\in{\mathbb{Z}}_{\geq 0}\text{ even}\\ &\\ -\dfrac{2^{-2n-1}}{\pi^{2n+2}}n!\,L_{0}(n+1)=-\dfrac{2^{-2n-1}}{\pi^{2n+2}}n!\,\displaystyle\sum_{m=1}^{\infty}\dfrac{A_{m}+A_{-m}}{m^{n+1}}&\text{ if }n\in{\mathbb{Z}}_{>0}\text{ odd}\end{cases}$ (73)
$=-\frac{2^{-2n-1}}{\pi^{2n+2}}n!\left(\sum_{m=1}^{\infty}\frac{A_{m}}{m^{n+1}}+\sum_{m=1}^{\infty}\frac{A_{-m}}{(-m)^{n+1}}\right)$
$=-\frac{2^{-2n-1}}{\pi^{2n+2}}n!\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{m^{n+1}}\,.$
To sum up, substituting Eq. (72) into Eq. (68) yields
$\tilde{f}_{+}(y)$ $=2\sum_{n=0}^{\infty}n!\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{(-4\pi^{2}m)^{n+1}}(2\pi{\rm i})^{n}y^{n}\,,$ (74)
$=\sum_{n=0}^{\infty}\left(\frac{n!}{\pi{\rm i}}\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{(2\pi{\rm i}m)^{n+1}}\right)y^{n}\,,$
which is a Gevrey-1 asymptotic series. Using Proposition 3.1, we conclude that the Borel transform of $y\tilde{f}_{+}(y)$ has a modular resurgent structure with simple poles at $\rho_{m}=2\pi{\rm i}m$, $m\in{\mathbb{Z}}_{\neq 0}$, and Stokes constants $2A_{m}$. Analogously, substituting Eq. (73) into Eq. (69) gives
$\tilde{f}_{-}(y)$ $=-2\sum_{n=0}^{\infty}n!\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{(4\pi^{2}m)^{n+1}}(2\pi{\rm i})^{n}y^{n}\,,$ (75)
$=-\sum_{n=0}^{\infty}\left(\frac{n!}{\pi{\rm i}}\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{-m}}{(2\pi{\rm i}m)^{n+1}}\right)y^{n}\,,$
which is a Gevrey-1 asymptotic series. Once more, using Proposition 3.1, we conclude that the Borel transform of $y\tilde{f}_{-}(y)$ has a modular resurgent structure with simple poles at $\rho_{m}=2\pi{\rm i}m$, $m\in{\mathbb{Z}}_{\neq 0}$, and Stokes constants $-2A_{-m}$. ∎
Additionally, the generating function defined in Eq. (66) satisfies Conjecture 2.
Proposition 4.2.
Within the assumptions of Proposition 4.1, the generating function $f:{\mathbb{C}}\setminus{\mathbb{R}}\to{\mathbb{C}}$ defined in Eq. (66) restricted to either the upper or the lower half of the complex plane is a weight-$1$ holomorphic quantum modular form for the group $\mathsf{SL}_{2}({\mathbb{Z}})$.
Proof.
Recall from part $(d)$ of Theorem $1$ in [31, Chapter 1] that there exists an analytic function $\psi\colon{\mathbb{C}}^{\prime}\to{\mathbb{C}}$ such that [^11] The function $\psi$ in Eq. (76) is the so-called period of the function $f$ in Eq. (66) and satisfies a three-terms functional equation. See [31] for details.
$c\,\psi(y)=f(y)-\frac{1}{y}f\left(-\frac{1}{y}\right)\,,$ (76)
for some constant $c\in{\mathbb{C}}_{\neq 0}$. Moreover, recall that the generators of the modular group $\mathsf{SL}_{2}({\mathbb{Z}})$ are the matrices
$T=\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\,,\quad S=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,.$ (77)
Hence, if we restrict the generating function to the upper half-plane, i.e., we take $f\colon{\mathbb{H}}\to{\mathbb{C}}$, the cocycle $h_{T}[f](y)$ is trivial due to the periodicity property $f(y+1)=f(y)$, while the cocycle for $S$ is given by
$h_{S}[f](y)=-c\,\psi(y)\,.$ (78)
The same arguments apply to the restriction of the generating function to the lower half-plane, i.e., for $f\colon{\mathbb{H}}_{-}\to{\mathbb{C}}$. ∎
With an extra assumption on the parity of the coefficients $A_{m}$, $m\in{\mathbb{Z}}_{\neq 0}$, appearing in the definition of the $L$-functions $L_{\epsilon}(s)$, $\epsilon=0,1$, in Eq. (63), we prove the effectiveness of the median resummation for the generating function defined in Eq. (66), which therefore satisfies Conjecture 1. In particular, assuming $A_{m}=A_{-m}$, we have that $f(-y)=-f(y)$ for $y\in{\mathbb{C}}\setminus{\mathbb{R}}$, so that $f(y)$ need only be specified in the upper half of the complex $y$-plane.
Proposition 4.3.
Within the assumptions of Proposition 4.1, if $A_{m}=A_{-m}$ for every $m\in{\mathbb{Z}}_{\neq 0}$, the generating function $f\colon{\mathbb{H}}\to{\mathbb{C}}$ defined in Eq. (66) is recovered from its asymptotic expansion $\tilde{f}_{+}(y)$ as $y\to 0$ with $\Im(y)>0$ through the median resummation. Precisely,
$yf(y)=\mathcal{S}^{\bf med}_{\frac{\pi}{2}}\big{[}y\tilde{f}_{+}(y)\big{]}\,,\quad y\in{\mathbb{H}}\,.$ (79)
Proof.
Due to the parity of the coefficients, $L_{1}(s)$ is identically zero, while
$L_{0}(s)=2\sum_{m=1}^{\infty}\frac{A_{m}}{m^{s}}\,,$ (80)
so that the two $L$-functions in Eqs. (67a) and (67b) reduce to
$L_{\pm}(s)=\frac{L_{0}(s)}{2}\,.$ (81)
As a consequence of Proposition 4.1, and particularly Eqs. (74) and (75), the formal power series $\tilde{f}_{\pm}(y)$ can be written as
$\tilde{f}_{\pm}(y)=\pm\frac{2}{\pi{\rm i}}\sum_{n=0}^{\infty}(2n+1)!\sum_{m>0}\frac{A_{m}}{\rho_{m}^{2n+2}}y^{2n+1}\,,$ (82)
where $\rho_{m}=2\pi{\rm i}m$ as before. Recall that the median resummation at angle $\pi/2$ can be expressed in terms of the Borel–Laplace sums at angles 00 and $\pi$ using Eq. (15). We have that
$\mathcal{S}^{\bf med}_{\frac{\pi}{2}}\big{[}y\tilde{f}_{+}(y)\big{]}$ =$=\begin{cases}s_{0}\big{(}y\tilde{f}_{+}(y)\big{)}+\frac{1}{2}\mathrm{disc}_{\frac{\pi}{2}}\big{[}y\tilde{f}_{+}(y)\big{]}\,,&\Re(y)>0\,,\\ \\ s_{\pi}\big{(}y\tilde{f}_{+}(y)\big{)}-\frac{1}{2}\mathrm{disc}_{\frac{\pi}{2}}\big{[}y\tilde{f}_{+}(y)\big{]}\,,&\Re(y)<0\,,\end{cases}$ (83)
for $\Im(y)>0$. To compute the Borel–Laplace sums of $\tilde{f}_{+}(y)$, let us first evaluate its Borel transform as in the proof of Proposition 3.1, that is,
$\mathcal{B}\big{[}y\tilde{f}_{+}(y)\big{]}(\zeta)=\frac{2}{\pi{\rm i}}\sum_{n=0}^{\infty}\sum_{m>0}\frac{A_{m}}{\rho_{m}^{2n+2}}\zeta^{2n+1}=-\frac{1}{\pi{\rm i}}\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{\zeta-\rho_{m}}\,,$ (84)
where we have exchanged the order of summation, resummed the geometric series over the index $n$, and applied the identity
$\frac{1}{(x-y)(x+y)}=\frac{1}{2x(x-y)}+\frac{1}{2x(x+y)}$ (85)
in the second step. Then, the Borel–Laplace sums at angles 00 and $\pi$ are
$s_{0}\big{(}y\tilde{f}_{+}(y)\big{)}$ $=-\frac{1}{\pi{\rm i}}\int_{0}^{\infty}{\rm e}^{-\zeta/y}\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{\zeta-\rho_{m}}\,d\zeta=-2\sum_{m\in{\mathbb{Z}}_{\neq 0}}A_{m}\mathbf{e}_{1}\left(-\frac{m}{y}\right)\,,\quad$ $y>0\,,$ (86a)
$s_{\pi}\big{(}y\tilde{f}_{+}(y)\big{)}$ $=-\frac{1}{\pi{\rm i}}\int_{0}^{-\infty}{\rm e}^{-\zeta/y}\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{\zeta-\rho_{m}}\,d\zeta=-2\sum_{m\in{\mathbb{Z}}_{\neq 0}}A_{m}\mathbf{e}_{1}\left(-\frac{m}{y}\right)\,,\quad$ $y<0\,,$ (86b)
where we have introduced the analytic function $\mathbf{e}_{1}\colon{\mathbb{C}}\setminus{\rm i}{\mathbb{R}}_{\geq 0}\to{\mathbb{C}}$ defined by [^12] The notation in Eq. (87) was suggested by M. Kontsevich as the function $\mathbf{e}_{1}$ plays a similar role in other examples (e.g., the case of $\sigma,\sigma^{*}$ in [23, from min. 21.55] and the spectral trace of local ${\mathbb{P}}^{2}$ in [19]).
$\mathbf{e}_{1}(z):=\frac{1}{2\pi{\rm i}}\int_{0}^{\infty}{\rm e}^{-2\pi t}\frac{dt}{t+{\rm i}z}=\frac{1}{2\pi{\rm i}}{\rm e}^{2\pi{\rm i}z}\,\Gamma(0,2\pi{\rm i}z)\,,$ (87)
where $\Gamma(s,2\pi{\rm i}z)$ denotes the upper incomplete gamma function. Note that the Borel–Laplace sums above can be analytically continued to $\Re(y)>0$ and $\Re(y)<0$, respectively, as a consequence of the properties of the function $\mathbf{e}_{1}$. Besides, by the standard residue argument, the discontinuity is
$\mathrm{disc}_{\frac{\pi}{2}}\big{[}y\tilde{f}_{+}(y)\big{]}=2\sum_{m>0}A_{m}{\rm e}^{-2\pi{\rm i}m/y}=2f\left(-\frac{1}{y}\right)\,,\quad\Im(y)>0\,.$ (88)
We can then apply the results of Section 3, Chapter II of [31], which uses the theory of even Maass cusp forms. In particular, adopting the notation of [31, Eqs. (2.11) and (2.13)], we define the function $\psi_{1}\colon{\mathbb{C}}\setminus{\rm i}{\mathbb{R}}_{\geq 0}\to{\mathbb{C}}$ as
$\psi_{1}(y):=2\pi{\rm i}\sum_{m\in{\mathbb{Z}}_{\neq 0}}A_{m}\mathbf{e}_{1}(my)\,,$ (89)
which satisfies the properties [31, Eq. (2.10)]
$y\psi_{1}(y)=\psi_{1}\left(\frac{1}{y}\right)\,,\quad\psi_{1}(-y)=\psi_{1}(y)\,,$ (90)
where the first formula holds for $\Re(y)>0$. Then, we have that [31, Eq. (2.20)] [^13] The same formula in Eq. (91) holds for $y\in{\mathbb{H}}_{-}$ with a change of sign in the constant $c_{*}$.
$c_{*}\,f(y)=\psi_{1}(y)+\frac{1}{y}\psi_{1}\left(-\frac{1}{y}\right)\,,$ (91)
for some constant $c_{*}\in{\mathbb{C}}_{\neq 0}$. Let us set the constant $c_{*}=-2\pi{\rm i}$. Using Eqs. (91) and (90), we deduce that
$yf(y)$ =$=\begin{cases}\displaystyle-\frac{1}{\pi{\rm i}}\psi_{1}\left(-\frac{1}{y}\right)+f\left(-\frac{1}{y}\right)\,,\quad&\Re(y)>0\,,\\ &\\ \displaystyle-\frac{1}{\pi{\rm i}}\psi_{1}\left(-\frac{1}{y}\right)-f\left(-\frac{1}{y}\right)\,,\quad&\Re(y)<0\,.\end{cases}$ (92)
It follows from Eqs. (89), (86a), and (86b) that we can identify $\psi_{1}(-1/y)$ with the Borel–Laplace sum of $y\tilde{f}_{+}(y)$ along the positive real axis for $\Re(y)>0$, that is,
$-\frac{1}{\pi{\rm i}}\psi_{1}\left(-\frac{1}{y}\right)=s_{0}\big{(}y\tilde{f}_{+}(y)\big{)}\,,\quad\Re(y)>0\,,$ (93)
and with its Borel–Laplace sum along the negative real axis for $\Re(y)<0$, that is,
$-\frac{1}{\pi{\rm i}}\psi_{1}\left(-\frac{1}{y}\right)=s_{\pi}\big{(}y\tilde{f}_{+}(y)\big{)}\,,\quad\Re(y)<0\,.$ (94)
Substituting the discontinuity in Eq. (88) and the Borel–Laplace sums in Eqs. (93) and (94) into the formula for the median resummation in Eq. (83), we find the same expression in Eq. (92). ∎
A result analogous to Proposition 4.3 applies for an odd sequence of complex numbers $A_{m}$, $m\in{\mathbb{Z}}_{\neq 0}$. In this case, the generating function in Eq. (66) satisfies $f(-y)=f(y)$ for $y\in{\mathbb{C}}\setminus{\mathbb{R}}$, while $L_{1}(s)=2\sum_{m=1}^{\infty}\tfrac{A_{m}}{m^{s}}$ is the only non-trivial $L$-function. Moreover, we expect a similar result for Maass cusps forms that are invariant under a discrete subgroup of $\mathsf{PSL}_{2}({\mathbb{Z}})$.
Remark 4.2.
The example of even Maass cusp forms discussed in Proposition 4.3 illustrates an interesting feature of modular resurgent series. Indeed, the cocycle $\psi(y)$ in Eq. (76) is proportional to the function $\psi_{1}(y)$ in Eq. (89) by means of Eq. (91). Consequently, it follows from Propositions 4.2 and 4.3 that the Borel–Laplace sums $s_{0}\big{(}y\tilde{f}_{\pm}(y)\big{)}$ are proportional to $\psi(-1/y)$. We expect this property to hold generally whenever both Conjectures 1 and 2 (or, equivalently, Conjectures 3 and 4) are verified. In other words, we expect the cocycles of $q$-series whose asymptotic expansions have a modular resurgent structure to appear naturally as their Borel–Laplace sums.
5 Conclusions
In this paper, we considered a particular class of resurgent asymptotic series that appear in numerous examples from the study of quantum invariants of knots and 3-manifolds, combinatorics, and topological string theory on toric CY threefolds. Motivated by the recent mathematical and physical interest, we introduced the notion of a modular resurgent series, whose Borel plane displays a single infinite tower of singularities, the secondary resurgent series are trivial, and the Stokes constants are coefficients of an $L$-function.
We proposed a new perspective on their modular resurgent structures by taking the point of view of the Stokes constants, their generating functions in the form of $q$-series, and their $L$-functions. Generalizing the strong-weak resurgent symmetry of the spectral trace of local ${\mathbb{P}}^{2}$ that we presented in the companion paper [19], we described the broader framework of modular resurgence. Pairs of modular resurgent structures are connected through a global exact symmetry revolving around the central role played by the functional equation satisfied by the $L$-functions whose coefficients are the Stokes constants. We illustrated the workings of modular resurgence via explicit computations in the concrete example of the generating functions of the Stokes constants for the weak and strong coupling asymptotic expansions of the spectral trace of local ${\mathbb{P}}^{2}$. Supported by disparate examples, we conjectured that the median resummation of modular resurgent series is effective and produces quantum modular forms. Finally, we proved that a large class of examples of modular resurgent series originating from the theory of Maass cusp forms satisfies our conjectural statements.
Among the possible questions one could address, we highlight two directions of research arising from the investigation performed in this paper. We would like to broaden the supporting basis for our paradigm of modular resurgence. A bigger pool of examples might help us deepen our understanding of the critical role of the $L$-functions constructed from the Stokes constants and their functional equation. A complete characterization of these ingredients might, in turn, steer us towards proving Conjectures 1 and 2.
Moreover, our statements on the median resummation could be placed in a more general context. Indeed, as the first author started addressing in [38], when divergent formal power series arise from the asymptotic expansion of well-defined analytic functions, it is natural to ask which summability method allows us to reconstruct the original functions. It would be interesting to study the effectiveness of summability methods in a broad sense, following the spirit of our Conjectures 1 and 3 and Conjecture 1.1 in [1].
Acknowledgements
We thank Alba Grassi, Maxim Kontsevich, Marcos Mariño, Campbell Wheeler, and Don Zagier for many useful discussions. This work has been supported by the ERC-SyG project “Recursive and Exact New Quantum Theory” (ReNewQuantum), which received funding from the European Research Council (ERC) within the European Union’s Horizon 2020 research and innovation program under Grant No. 810573 and by the Swiss National Centre of Competence in Research SwissMAP (NCCR 51NF40-141869 The Mathematics of Physics).