### v2

# Modular resurgent structures

Abstract
The theory of resurgence uniquely associates a factorially divergent formal power series with a collection of exponentially small non-perturbative corrections paired with a set of complex numbers known as Stokes constants. When the Borel plane displays a single infinite tower of singularities, the secondary resurgent series are trivial, and the Stokes constants are coefficients of an $L$-function, a rich analytic number-theoretic fabric underlies the resurgent structure of the asymptotic series. We propose a new paradigm of modular resurgence that focuses on the role of the Stokes constants and the interplay of the $q$-series acting as their generating functions with the corresponding $L$-functions. Guided by two pivotal examples arising from topological string theory and the theory of Maass cusp forms, we introduce the notion of modular resurgent series, which we conjecture to have specific summability properties as well as to be related to quantum modular forms.
1 Introduction
Resurgence, summability, and quantum modularity are very active domains of research involving different communities, which have so far mainly focused on the study of various examples coming from quantum Chern–Simons theory, in particular from the quantum invariants of knots and 3-manifolds [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. There is currently no general theory that integrates these three research areas into a comprehensive framework. In this paper, we aim to address this gap.
We take advantage of the abundant amount of information gathered from the in-depth analysis of the resurgence of the spectral trace of the toric Calabi–Yau (CY) threefold known as local ${\mathbb{P}}^{2}$ performed by the second author in [19]. This satisfies a full-fledged strong-weak resurgent symmetry and special modularity and summability properties that we present in the companion paper [20]. Generalizing the cardinal features of this example, we characterize those resurgent asymptotic series that produce holomorphic quantum modular forms and, at the same time, whose median resummation is effective. This newly proposed paradigm goes under the name of modular resurgence and deeply relies on the fundamental role played by the Stokes constants and the $q$-series acting as their generating functions. Our results naturally lead us to present the definition of a modular resurgent series, whose Borel plane displays a single infinite tower of singularities, the secondary resurgent series are trivial, and the Stokes constants are coefficients of an $L$-function. The resurgent structure of such a series is called a modular resurgent structure.
We show that a global network of exact relations connects the complete resurgent structures of pairs of modular resurgent series, which is built upon the analytic properties of the $L$-functions whose coefficients are the Stokes constants, as illustrated in the diagram in Eq. (41). Guided by a diverse and conspicuous set of examples that provide supporting evidence, we propose in Conjectures 1 and 2 that those $q$-series whose perturbative series is modular resurgent are holomorphic quantum modular forms and are reconstructed by the median resummation of their asymptotic expansions. Beyond the complete, original example of the spectral trace of local ${\mathbb{P}}^{2}$ of [19, 20], which is briefly discussed in Section 4.1, our statements are proven in Theorems 4.3 and 4.2 for a large class of modular resurgent series originating from the theory of Maass cusp forms. Finally, we build in Theorem 4.4 another class of supporting examples built from cusp forms. More generally, evidence of our conjectures comes from studying the quantum invariants of knots and 3-manifolds, combinatorics, and topological strings on toric CY threefolds.
Quantum invariants.
The example of the trefoil knot and other families of toric knots have been discussed separately from the resurgence, summability, and modularity perspectives. Evidence in support of our conjectures is present in the works of [1, 21]. Further evidence arises from the study of 3-manifold invariants, such as the Witten–Reshetikhin–Turaev (WRT) invariants. The authors of [2] prove that the Ohtsuki series $Z_{0}$ for Seifert fibered homology spheres satisfies the conditions of what we define here as a modular resurgent series. Its median resummation reconstructs the normalized WRT invariants $\tilde{Z}_{0}$, which are related to the Gukov–Pei–Putrov–Vafa (GPPV) invariants $\hat{Z}_{0}$ of [22]. Then, the results of [13, 14] show that the GPPV invariants of certain plumbed 3-manifolds are quantum modular forms.
Combinatorics.
Families of $q$-hypergeometric series [23] provide instances of quantum modular forms, among which the example of $\sigma,\sigma^{*}$ first discussed by Zagier in [6, Example 1]. The resurgent structure of $\sigma,\sigma^{*}$ [24] has been studied by the first author and M. Kontsevich and is modular. Furthermore, we prove here that a larger class of examples built from Maass cusp forms with spectral parameter $1/2$ and more general cusp forms satisfy our conjectures.
Topological strings.
In the context of the TS/ST correspondence [25, 26], the only example we are aware of comes from the results obtained in [19, 20] on the resurgence of the spectral trace of local ${\mathbb{P}}^{2}$ at strong and weak coupling. However, we expect the existence of an appropriate generalization to other toric CY threefolds and higher-order fermionic spectral traces.
This paper is organized as follows. In Section 2, we review the basic ingredients in the resurgent analysis of formal power series and the definitions of (holomorphic) quantum modular forms. In Section 3, we delve into the analytic number-theoretic characterization of those resurgent structures that give rise to quantum modular forms according to the newly introduced paradigm of modular resurgence. We present the definition of a modular resurgent series, which we conjecture to have peculiar modularity and summability properties. In Section 4, we review the example of the spectral trace of local ${\mathbb{P}}^{2}$, briefly summarizing the strong-weak resurgent symmetry of [20], and prove our conjectures for a set of examples originating from the theory of Maass cusp forms and more general modular forms. Finally, in Section 5, we conclude and mention further perspectives to be addressed in future works.
2 Basics of resurgence and modularity
In this section, we review the necessary background notions in resurgence and quantum modularity. The resurgent analysis of factorially divergent formal power series unveils a universal mathematical structure of non-perturbative sectors involving a collection of numerical data called Stokes constants [27]. See [28, 29, 30, 31] for an introduction to the theory of resurgence.
Modular forms have a long history and are of interest in different fields. Generally speaking, they outline the invariant properties of certain $q$-series under the action of the modular group. However, there are examples of $q$-series that are not modular forms and yet whose failure of modularity can be described explicitly. These belong to the class of quantum modular forms introduced by Zagier in [6]. We refer to [32, Chapter IV] and references therein for an overview.
2.1 Resurgent asymptotic series
The Borel transform ${\cal B}\colon z^{-\alpha}{\mathbb{C}}[\![z]\!]\to\zeta^{-\alpha}{\mathbb{C}}\{\zeta\}$, where $z,\zeta$ are formal variables and $\alpha\in{\mathbb{R}}\backslash{\mathbb{Z}}_{\geq 0}$, acts on $z$-monomials as
$\mathcal{B}[z^{n-\alpha}]:=\frac{\zeta^{n-\alpha-1}}{\Gamma(n-\alpha)}\,,\quad n\in{\mathbb{Z}}_{\geq 0}\,,$ (1)
where $\Gamma(n-\alpha)$ is the gamma function. The action of the Borel transform above is extended by linearity to all formal power series in $z^{-\alpha}{\mathbb{C}}[\![z]\!]$, while the Borel transform of the identity is conventionally denoted with the symbol $\delta:=\mathcal{B}[1]$. Let $\phi(z)$ be a Gevrey-1 asymptotic series of the form
$\phi(z)=z^{-\alpha}\sum_{n=0}^{\infty}a_{n}z^{n}\in z^{-\alpha}{\mathbb{C}}[\![z]\!]\,,\quad|a_{n}|\leq{\cal A}^{-n}n!\quad n\gg 1\,,$ (2)
where $\alpha\in{\mathbb{R}}\backslash{\mathbb{Z}}_{\geq 0}$ and ${\cal A}\in{\mathbb{R}}_{>0}$. Its Borel transform, which we denote for simplicity by $\hat{\phi}(\zeta):=\mathcal{B}\phi(\zeta)$, is defined as
$\hat{\phi}(\zeta)=\sum_{k=0}^{\infty}\frac{a_{k}}{\Gamma(k-\alpha)}\zeta^{k-\alpha-1}\in\zeta^{-\alpha}{\mathbb{C}}\{\zeta\}\,,$ (3)
which is holomorphic in the open disk centered at $\zeta=0$ of radius ${\cal A}$. When extended to the whole complex $\zeta$-plane, known as the Borel plane, $\hat{\phi}(\zeta)$ shows a (possibly infinite) set of singularities $\zeta_{\omega}\in{\mathbb{C}}$, which we label by the index $\omega\in\Omega$. A ray in the Borel plane of the form
${\cal C}_{\theta_{\omega}}={\rm e}^{{\rm i}\theta_{\omega}}{\mathbb{R}}_{\geq 0}\,,\quad\theta_{\omega}=\arg(\zeta_{\omega})\,,$ (4)
which starts at the origin and passes through a singularity $\zeta_{\omega}$, is called a Stokes ray. The Borel plane is partitioned into sectors bounded by the Stokes rays in such a way that the Borel transform converges to a generally different holomorphic function in each of these sectors.
Definition 2.1.
A Gevrey-1 series $\phi\in z^{-\alpha}{\mathbb{C}}[\![z]\!]_{1}$ is resurgent if its Borel transform $\hat{\phi}\in\zeta^{-\alpha}{\mathbb{C}}\{\zeta\}$ can be endlessly analytically continued [27]. Namely, for every $L>0$, there is a finite set of points $\Omega_{L}$ in the Riemann surface of $\zeta^{-\alpha}$ such that $\hat{\phi}$ can be analytically continued along any path of length at most $L$ that avoids $\Omega_{L}$. If, additionally, the Borel transform of a resurgent series $\phi$ has only simple poles and logarithmic branch points, it is called simple resurgent.
Let us assume that the formal power series in Eq. (2) is simple resurgent. If the singularity $\zeta_{\omega}$ is a simple pole, the local expansion around $\zeta=\zeta_{\omega}$ of the Borel transform in Eq. (3) has the form
$\hat{\phi}(\zeta)=-\frac{S_{\omega}}{2\pi{\rm i}(\zeta-\zeta_{\omega})}+\text{regular in $\zeta-\zeta_{\omega}$}\,,$ (5)
where $S_{\omega}\in{\mathbb{C}}$ is the Stokes constant at $\zeta_{\omega}$. Whereas, if the singularity $\zeta_{\omega}$ is a logarithmic branch point, the local expansion around $\zeta=\zeta_{\omega}$ of the Borel transform is
$\hat{\phi}(\zeta)=-\frac{S_{\omega}}{2\pi{\rm i}}\log(\zeta-\zeta_{\omega})\hat{\phi}_{\omega}(\zeta-\zeta_{\omega})+\text{regular in $\zeta-\zeta_{\omega}$}\,,$ (6)
where again $S_{\omega}\in{\mathbb{C}}$ is the Stokes constant and the formal series
$\hat{\phi}_{\omega}(\zeta)=\sum_{k=0}^{\infty}\hat{a}_{k,\omega}\zeta^{k-\beta}\in\zeta^{-\beta}{\mathbb{C}}\{\zeta\}\,,$ (7)
where $\beta\in{\mathbb{R}}\setminus{\mathbb{Z}}_{\geq 0}$, can be regarded as the Borel transform of the Gevrey-1 asymptotic series
$\phi_{\omega}(z)=z^{-\beta}\sum_{n=0}^{\infty}a_{n,\omega}z^{n}\in z^{-\beta}{\mathbb{C}}[\![z]\!],\,\quad a_{n,\omega}=\Gamma(n-\beta+1)\,\hat{a}_{n,\omega}\,.$ (8)
Note that the value of the Stokes constant $S_{\omega}$ depends on the normalization of $\phi_{\omega}$
Remark 2.1.
One of the fundamental characteristics of resurgent asymptotic series is the emergence of new series from the behavior of the analytic continuation of their Borel transform near a singularity. More precisely, in the case of a simple pole, the Stokes constant given by the residue captures the new information, namely, $S_{\omega}$ in Eq. (5). When the singularity is a logarithmic branch point, a new germ of an analytic function resurges from the local behavior of the Borel transform, namely, $S_{\omega}\hat{\phi}_{\omega}$ in Eq. (6).
If the analytic continuation of the Borel transform in Eq. (3) does not grow too fast at infinity at an angle $0\leq\theta<2\pi$ in the Borel plane, [^1] We require that $\hat{\phi}(\zeta)$ grows at most exponentially in a tubular neighbourhood of the Borel plane containing the angle $\theta$ [29, Section 5.6]. its Laplace transform along the direction $\theta$ gives the Borel–Laplace sum of the original, divergent formal power series $\phi(z)$ in the same direction, which we denote by $s_{\theta}(\phi)(z)$. Explicitly,
[…]
whose asymptotics near the origin reconstructs $\phi(z)$ [29, Theorem 5.20].
Definition 2.2.
Let $\phi\in z^{-\alpha}{\mathbb{C}}[\![z]\!]$ be a Gevrey-1 resurgent asymptotic series. If the Borel–Laplace sum of $\phi$ at an angle $0\leq\theta<2\pi$ as in Eq. (9) exists in some region of the complex $z$-plane, we say that the series $\phi$ is Borel–Laplace summable along the direction $\theta$.
Note that the Borel–Laplace sum inherits the sectorial behavior of the Borel transform. Specifically, it is a locally analytic function with discontinuities across the special rays identified by
$\arg(z)=\arg(\zeta_{\omega})\,,\quad\omega\in\Omega\,.$ (10)
Definition 2.3.
The discontinuity across an arbitrary ray ${\cal C}_{\theta}={\rm e}^{{\rm i}\theta}{\mathbb{R}}_{\geq 0}$ in the complex $z$-plane is the difference between the Borel–Laplace sums along two rays that lie slightly above and slightly below ${\cal C}_{\theta}$. Namely,
[…]
where $\theta_{\pm}=\theta\pm\epsilon$ for some small positive angle $\epsilon$.
A standard contour deformation argument shows that the two lateral Borel–Laplace sums differ by exponentially small terms. More precisely, if the Borel transform $\hat{\phi}(z)$ has only simple poles, then we have that
$\mathrm{disc}_{\theta}\phi(z)=\sum_{\omega\in\Omega_{\theta}}S_{\omega}{\rm e}^{-\zeta_{\omega}/z}\,,$ (12)
where the index $\omega\in\Omega_{\theta}$ labels the singularities $\zeta_{\omega}$ such that $\arg(\zeta_{\omega})=\theta$ and the complex numbers $S_{\omega}$ are the same Stokes constants that appear in Eq. (5). Similarly, if there are only logarithmic branch points, we have that
$\mathrm{disc}_{\theta}\phi(z)=\sum_{\omega\in\Omega_{\theta}}S_{\omega}{\rm e}^{-\zeta_{\omega}/z}s_{\theta_{-}}(\phi_{\omega})(z)\,,$ (13)
where again the complex numbers $S_{\omega}$ are the Stokes constants appearing in Eq. (6), while $\phi_{\omega}(z)$ is the formal power series in Eq. (8).
The effectiveness of the median resummation
In the study of the summability properties of factorially divergent formal power series that arise as asymptotic limits of well-defined holomorphic functions, it is natural to ask which summability method is efficient in reproducing the original function from its asymptotic expansion. In some examples, like the ones discussed in [33], the Borel–Laplace sum is effective. However, to reconstruct certain $q$-series by resumming their asymptotic expansion, sometimes one has to consider the so-called median resummation.
Definition 2.4.
The median resummation of the Gevrey-1 asymptotic series $\phi(z)\in{\mathbb{C}}\llbracket z\rrbracket$ in Eq. (2) across an arbitrary ray ${\cal C}_{\theta}={\rm e}^{{\rm i}\theta}{\mathbb{R}}_{\geq 0}$ in the complex $z$-plane is defined as the average of the two lateral Borel–Laplace sums $s_{\theta_{\pm}}(\phi)(z)$, that is,
$\mathcal{S}^{\mathrm{med}}_{\theta}\phi(z):=\frac{s_{\theta_{+}}(\phi)(z)+s_{\theta_{-}}(\phi)(z)}{2}\,.$ (14)
Notice that $\mathcal{S}^{\mathrm{med}}_{\theta}\phi(z)$ is an analytic function for $\arg z\in(\theta-\tfrac{\pi}{2},\theta+\tfrac{\pi}{2})$. In addition, by a contour deformation argument, we can equivalently write it as
$\mathcal{S}^{\mathrm{med}}_{\theta}\phi(z)=\begin{cases}s_{\theta_{-}}(\phi)(z)+\frac{1}{2}\,\mathrm{disc}_{\theta}\phi(z)\,,\quad&\Re\left({\rm e}^{-{\rm i}\theta_{-}}z\right)>0\,,\\ &\\ s_{\theta_{+}}(\phi)(z)-\frac{1}{2}\,\mathrm{disc}_{\theta}\phi(z)\,,\quad&\Re\left({\rm e}^{-{\rm i}\theta_{+}}z\right)>0\,,\end{cases}$ (15)
where $\Re\left({\rm e}^{-{\rm i}\theta}z\right)>0$. The expression in Eq. (15) highlights how the discontinuities can be suitably interpreted as corrections to the standard Borel–Laplace sums $s_{\theta_{\pm}}(\phi)(z)$. In addition, by varying the contour of integration of the Laplace transform, the domain of analyticity of the median resummation can be extended.
In the context of quantum Chern–Simons theory, Costin and Garoufalidis conjectured that the asymptotic expansions at roots of unity of quantum invariants, such as the Kashaev invariant for knots and the WRT invariant for $3$-manifolds, are Gevrey-1 asymptotic series whose median resummation reconstructs the non-perturbative invariants [1, Conjecture 1]. As we show in [20, Theorem 4.6], further evidence of the efficacy of the median resummation comes from the study of the spectral traces of quantum-mechanical operators obtained by quantizing the mirror curve to a toric CY threefold.
2.2 Quantum modular forms
A first description of quantum modular forms was given by Zagier in [6]. Namely, a function $f\colon{\mathbb{Q}}\to{\mathbb{C}}$ is called a weight-$\omega$ quantum modular form with respect to a subgroup $\Gamma\subseteq\mathsf{SL}_{2}({\mathbb{Z}})$, where $\omega$ is a fixed integer or half-integer, if the cocycle
$h_{\gamma}[f](y):=(cy+d)^{-\omega}f\left(\frac{ay+b}{cy+d}\right)-f(y)$ (16)
is better behaved than $f(y)$ for all choices of $\gamma=\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\in\Gamma$. More precisely, being better behaved means having better analyticity properties than the function $f(y)$ itself—e.g., being a real analytic function over ${\mathbb{R}}\setminus\{\gamma^{-1}(\infty)\}$. Quantum modular forms of weight zero are called quantum modular functions.
Some examples of quantum modular forms are built from modular forms. For instance, it was proven in [21] that the function $g\colon{\mathbb{Q}}\to{\mathbb{C}}$ defined by
$g(y):=q^{1/24}\Big{(}1+\sum_{n=1}^{\infty}(1-q)(1-q^{2})\ldots(1-q^{n})\Big{)}\,,\quad q={\rm e}^{2\pi{\rm i}y}\,,$ (17)
which is closely related to the Kashaev invariant for the trefoil knot [1] [^2] Other examples arising from the quantum invariants of hyperbolic knots have been discussed in the literature, although their cocycles generally have a more complicated structure [7]. It is unknown whether they are related to modular forms., is a weight-$3/2$ quantum modular form for $\mathsf{SL}_{2}({\mathbb{Z}})$. Its cocycle $h_{\gamma}[g]$ is written explicitly as
[…]
where $\eta(y)=\sum_{k=0}^{\infty}\chi(k)q^{k}$ is the Dedekind eta function and $\chi$ is the Dirichlet character modulo $12$ with $\chi(\pm 1)=1$ and $\chi(\pm 5)=-1$. As shown in [21], the function $g$ agrees to all orders with the radial limit of the Eichler integral of $\eta$, that is,
$g(y)``="-\frac{1}{2}\sum_{k=0}^{\infty}k\chi(k)q^{k}\,,\quad y\in{\mathbb{Q}}\,.$ (19)
Besides, the cocycle in Eq. (18) is real analytic on ${\mathbb{R}}\setminus\{\gamma^{-1}(\infty)\}$ and is expressed in terms of the original weight-1/2 modular form $\eta$. This is a common feature of those quantum modular forms that arise from a classical counterpart, and it is manifest in other examples. See [6, Examples 1, 3, 4] and [23, 12, 13, 14]. Other examples from the quantum invariants of hyperbolic knots have cocycles with a more complicated structure [7], and whether they are related to modular forms is unknown.
Notice that the Eichler integral on the right-hand side (RHS) of Eq. (19) is well-defined and holomorphic in the upper half of the complex plane, which we denote by ${\mathbb{H}}$. Indeed, it is also natural to ask whether holomorphic functions on ${\mathbb{H}}$ can be quantum modular forms. Although a holomorphic function $f\colon{\mathbb{H}}\to{\mathbb{C}}$ already has good analyticity properties, a version of quantum modularity is obtained by requiring that the cocycle $h_{\gamma}[f]$ in Eq. (16) is analytic in a domain larger than ${\mathbb{H}}$. The quantum modular forms that belong to this class of examples are called holomorphic quantum modular forms. The following definition was proposed by Zagier [34, min 21:45].
Definition 2.5.
An analytic function $f\colon{\mathbb{H}}\to{\mathbb{C}}$ is a weight-$\omega$ holomorphic quantum modular form for a subgroup $\Gamma\subseteq\mathsf{SL}_{2}({\mathbb{Z}})$, where $\omega$ is integer or half-integer, if the cocycle $h_{\gamma}[f]\colon{\mathbb{H}}\to{\mathbb{C}}$ in Eq. (16) extends holomorphically to [^3] Note that ${\mathbb{C}}_{\gamma}={\mathbb{C}}\setminus\left(-\infty;\,-d/c\right]$ when $c>0$ and ${\mathbb{C}}_{\gamma}={\mathbb{C}}\setminus\left[-d/c;\,+\infty\right)$ when $c<0$.
${\mathbb{C}}_{\gamma}:=\{y\in{\mathbb{C}}\colon cy+d\in{\mathbb{C}}\setminus{\mathbb{R}}_{\leq 0}\}$ (20)
for every $\gamma=\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\in\Gamma$.
We will also adopt the standard notation ${\mathbb{C}}^{\prime}={\mathbb{C}}\setminus{\mathbb{R}}_{\leq 0}$. For instance, following Example 2 in [6], let $\Delta\colon{\mathbb{H}}\to{\mathbb{C}}$ be the modular discriminant of the Dedekind eta function $\eta$, which is given by
$\Delta(y):=\eta(y)^{24}=q\prod_{n=1}^{\infty}(1-q^{n})^{24}\,,\quad q={\rm e}^{2\pi{\rm i}y}\,.$ (21)
It was shown in [6, Example 2] that its Eichler integral $\tilde{\Delta}\colon{\mathbb{H}}\to{\mathbb{C}}$, that is,
[…]
is a weight-10 holomorphic quantum modular form for $\mathsf{SL}_{2}({\mathbb{Z}})$, whose cocycle is known as
[…]
In particular, the cocycle in Eq. (23) is a polynomial of degree less than or equal to $10$ and therefore holomorphic on ${\mathbb{C}}$. This is an exceptional example since, typically, the cocycle only extends to some cut plane, e.g., ${\mathbb{C}}\setminus{\mathbb{R}}_{\leq 0}$. In addition, $h_{\gamma}[\tilde{\Delta}]$ has a convergent power series expansion as $y\to 0$. It follows that our resurgence proposal cannot be applied to this particular example. See also Remark 3.2.
Other examples of holomorphic quantum modular forms can be constructed from the theory of Maass cusp forms and their periods; see Section 4.2. Moreover, and remarkably, as noticed in [20, Section 4.3], the resurgence of the spectral trace of local ${\mathbb{P}}^{2}$ is a new source of examples of holomorphic quantum modular forms; see Section 4.1. Note that an analog of Definition 2.5 can be written for a function $f\colon{\mathbb{H}}_{-}\to{\mathbb{C}}$, where ${\mathbb{H}}_{-}$ denotes the lower half of the complex plane. Even more generally, a notion of vector-valued quantum modular forms on ${\mathbb{H}}\cup{\mathbb{Q}}\cup{\mathbb{H}}_{-}$ was proposed in [32, Definition 22].
We end this section with two remarks. Firstly, the adjective quantum referred to modular forms indicates the failure of invariance by the action of the modular group $\Gamma\subseteq\mathsf{SL}_{2}({\mathbb{Z}})$. For example, the notion of quantum Jacobi forms was introduced in [32, Definition 23] to describe the quantum modularity properties of certain $q$-Pochhammer symbols discussed in [32, Section 8.11, Theorem A-41]. Secondly, quantum modular forms constitute a much larger class than modular forms, and characterizing them might look rather complicated. However, when the failure of modularity is measured by a better-behaved analytic function with divergent asymptotics, we propose using the resurgence theory to identify a common structure beyond the divergent asymptotic behavior of the cocycle. This argument is detailed and discussed in Section 3.2.
3 Modular resurgence
In this section, we consider a class of Gevrey-1 asymptotic series that replicate the central features of the resurgence of the spectral trace of local ${\mathbb{P}}^{2}$, which has been extensively studied in [19, 20]. We propose a new perspective on their resurgent structures by taking the point of view of the $q$-series acting as generating functions of the Stokes constants. In particular, a cardinal role in this paradigm of modular resurgence is played by the $L$-functions whose coefficients are given by the Stokes constants and their functional equation. We describe a global exact resurgent symmetry connecting pairs of what we define as modular resurgent series and present a conjectural relation between the latter and quantum modular forms.
3.1 New perspectives on resurgence via $L$-functions
Recall that one of the main features of resurgent series is that one can compute the non-perturbative information hidden in the asymptotic expansion by studying the structure of singularities in the Borel plane. Namely, starting from the germ at the origin, that is, the Borel transform of the divergent formal power series, and analytically continuing it to the whole complex plane, it is possible to recover the germs at the other singularities, as discussed in Section 2.1. In the example of the spectral trace of local ${\mathbb{P}}^{2}$, the resurgent structure appears trivial at first glance, as the germs at the other singularities are all constants. Yet, as discovered in [19] and fully outlined in [20], when regarded as a sequence, these constants possess rich analytic number-theoretic properties that make contact with the notion of quantum modularity. Furthermore, when considering both weak and strong coupling limits, a unique strong-weak resurgent symmetry holds. Since these features can be generalized to a whole new class of asymptotic series, the results of [20] encourage us to adopt a novel perspective and motivate the following discussion.
We will consider a particular type of resurgent series and present a new paradigm of resurgence built upon the Stokes constants, their generating functions, and their Dirichlet series. We refer to this new paradigm as modular resurgence for reasons that will become clear in Section 3.2. Let us start by formally presenting and proving all relevant statements. The following result is a generalization of the exact large-order relations of [19, Sections 4.2.3 and 4.3.3].
Proposition 3.1.
Let $\tilde{f}(y)=\sum_{n=1}^{\infty}c_{n}y^{n}\in{\mathbb{C}}[\![y]\!]$ be a Gevrey-1 asymptotic series. If its coefficients are given explicitly by
$c_{n}=\frac{\Gamma(n)}{2\pi{\rm i}}\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{\zeta_{m}^{n}}\,,\quad n\in{\mathbb{Z}}_{>0}\,,$ (24)
where $A_{m}$, $\zeta_{m}$, $m\in{\mathbb{Z}}_{\neq 0}$, are complex numbers, then its Borel transform is a resurgent function with only simple poles at the locations $\zeta_{m}$ and corresponding Stokes constants $A_{m}$.
Proof.
By definition, the Borel transform of $\tilde{f}$ is the function
${\cal B}[\tilde{f}](\zeta)=\sum_{n=1}^{\infty}\frac{c_{n}}{\Gamma(n)}\zeta^{n-1}=\frac{1}{2\pi{\rm i}}\sum_{n=1}^{\infty}\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{\zeta_{m}^{n}}\zeta^{n-1}\in{\mathbb{C}}\{\zeta\}\,.$ (25)
Since $\tilde{f}$ is Gevrey-1, the coefficients $c_{n}$ have the factorially divergent growth
$|c_{n}|\leq n!|{\cal E}|^{-n}\quad n\gg 1\,,\quad{\cal E}\in{\mathbb{C}}\,.$ (26)
Hence, we can change the order of summation in Eq. (25), obtaining
${\cal B}[\tilde{f}](\zeta)=\frac{1}{2\pi{\rm i}}\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{\zeta_{m}}\sum_{n=1}^{\infty}\left(\frac{\zeta}{\zeta_{m}}\right)^{n-1}=-\frac{1}{2\pi{\rm i}}\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{\zeta-\zeta_{m}}\,,$ (27)
which has simple poles at $\zeta=\zeta_{m}$ with corresponding Stokes constants $A_{m}$. ∎
Note that Eq. (24) can be equivalently written as
$c_{n}=\frac{\Gamma(n)}{(2\pi{\rm i}){\cal E}^{n}}\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{m^{n}}\,,\quad n\in{\mathbb{Z}}_{>0}\,,$ (28)
when $\zeta_{m}={\cal E}m$, $m\in{\mathbb{Z}}_{\neq 0}$, for some constant ${\cal E}\in{\mathbb{C}}$. In addition, when the Stokes constants have moderate growth, the converse of the previous statement holds. This applies to the case of Stokes constants that are the coefficients of an $L$-function.
Proposition 3.2.
Let $\tilde{f}(y)=\sum_{n=1}^{\infty}c_{n}y^{n}\in{\mathbb{C}}[\![y]\!]$ be a Gevrey-1 asymptotic series. If its Borel transform is a resurgent function with only simple poles at the locations $\zeta_{m}\in{\mathbb{C}}$, $m\in{\mathbb{Z}}_{\neq 0}$, and the corresponding Stokes constants $A_{m}\in{\mathbb{C}}$ have moderate growth as $|m|\gg 1$, then the coefficients $c_{n}$ are given by Eq. (24).
Proof.
If the Borel transform of $\tilde{f}$ has simple poles at $\zeta_{m}$ with residues $A_{m}$ for $m\in{\mathbb{Z}}_{\neq 0}$, then it has the form
${\cal B}[\tilde{f}](\zeta)=-\frac{1}{2\pi{\rm i}}\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{\zeta-\zeta_{m}}\,,$ (29)
which is absolutely convergent thanks to the assumption on the growth of the Stokes constants $A_{m}$. We can then expand the geometric series and exchange the summation order, which yields
${\cal B}[\tilde{f}](\zeta)=\frac{1}{2\pi{\rm i}}\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{\zeta_{m}}\sum_{n=0}^{\infty}\left(\frac{\zeta}{\zeta_{m}}\right)^{n}=\frac{1}{2\pi{\rm i}}\sum_{n=1}^{\infty}\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{\zeta_{m}^{n}}\zeta^{n-1}\,.$ (30)
Then, by applying the definition of the Borel transform in Eq. (1), we obtain the expected expression for $c_{n}$. ∎
Up to the prefactors, the RHS in Eq. (28) is the Dirichlet series with coefficients given by the Stokes constants $A_{m}$ evaluated at the positive integer $n$. This does not a priori define an $L$-series. However, when it does, we can say something more.
Proposition 3.3.
Let $\{A_{m}\}$, $m\in{\mathbb{Z}}_{\neq 0}$, be a sequence of complex numbers defining the $L$-series
$L_{+}(s)=\sum_{m>0}\frac{A_{m}}{m^{s}}\,,\quad L_{-}(s)=-\sum_{m>0}\frac{A_{-m}}{m^{s}}$ (31)
that are convergent in some right half-plane $\{\Re(s)>\alpha\geq 0\}\subset{\mathbb{C}}$. Let
$f(y)=\begin{cases}\displaystyle\sum_{m>0}A_{m}{\rm e}^{2\pi{\rm i}my}&\mbox{if}\quad\Im(y)>0\\ &\\ -\displaystyle\sum_{m<0}A_{m}{\rm e}^{2\pi{\rm i}my}&\mbox{if}\quad\Im(y)<0\end{cases}$ (32)
be the generating series of the complex numbers $A_{m}$, $m\in{\mathbb{Z}}_{\neq 0}$. Then, $f(y)$ converges for $y\in{\mathbb{C}}\setminus{\mathbb{R}}$, and there is a canonical correspondence between the pair of $L$-series $L_{+}(s),\,L_{-}(s)$ and the generating function $f(y)$.
Proof.
The statement follows from the properties of the Mellin transform. Indeed, a simple computation shows that
[…]
that is, the $L$-series $L_{\pm}(s)$ are given by the Mellin transform of the generating function $f(y)$. Conversely, by the Mellin inversion theorem, $f(y)$ can be computed via the inverse Mellin transform of $L_{\pm}(s)$. Indeed, for $C>\alpha$ fixed, we have that
[…]
which can be analytically continued to all arguments $y\in{\mathbb{C}}\setminus{\mathbb{R}}$. ∎
Let us stress that the generating function $f(y)$ in Eq. (32) has the form of a $q$-series for $q={\rm e}^{2\pi{\rm i}y}$ when restricted to either the upper or the lower halves of the complex plane. Besides, we could equivalently consider another pair of $L$-series, that is,
$L_{0}(s)=\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{|m|^{s}}\,,\quad L_{1}(s)=\sum_{m\in{\mathbb{Z}}_{\neq 0}}\mathrm{sgn}(m)\frac{A_{m}}{|m|^{s}}\,,$ (35)
which are again in a one-to-one correspondence with the generating function $f(y)$ in Eq. (32) as a consequence of Proposition 3.3. Indeed, $L_{0}(s)$ and $L_{1}(s)$ can be written as the linear combinations
$L_{0}(s)$ $=\sum_{m>0}\frac{A_{m}}{m^{s}}+\sum_{m<0}\frac{A_{m}}{(-m)^{s}}=L_{+}(s)-L_{-}(s)\,,$ (36)
$L_{1}(s)$ $=\sum_{m>0}\frac{A_{m}}{m^{s}}-\sum_{m<0}\frac{A_{m}}{(-m)^{s}}=L_{+}(s)+L_{-}(s)\,.$
We can then consider the meromorphic extension of these $L$-series, or the $L$-series in Eq. (31), to the whole complex $s$-plane, that is, the corresponding $L$-functions. The following result states that knowing in closed form the perturbative coefficients of the asymptotic series obtained by expanding the generating function $f(y)$ at zero for $\Im(y)>0$ is equivalent to computing the analytic continuation of the $L$-series $L_{+}(s)$. An analogous statement holds for $\Im(y)<0$.
Proposition 3.4.
Let $\{A_{m}\}$, $m\in{\mathbb{Z}}_{>0}$, be a sequence of complex numbers defining an $L$-series
$L(s)=\sum_{m=1}^{\infty}\frac{A_{m}}{m^{s}}$ (37)
that is convergent in some right half-plane $\{\Re(s)>\alpha\geq 0\}\subset{\mathbb{C}}$ and can be analytically continued to $\{\Re(s)<0\}\subset{\mathbb{C}}$ through a functional equation. Let
$\tilde{f}(y)=\sum_{n=0}^{\infty}c_{n}y^{n}\in{\mathbb{C}}[\![y]\!]$ (38)
be the asymptotic series of the generating function $f(y)$, defined in Eq. (32), as $y\to 0$ with $\Im(y)>0$. Then, the perturbative coefficients $c_{n}$, $n\in{\mathbb{Z}}_{\geq 0}$, are given by
$c_{n}=L(-n)\frac{(2\pi{\rm i})^{n}}{n!}\,,$ (39)
where $L(-n)$ is defined by the functional equation.
Proof.
Taking the formal power series expansion of the exponential ${\rm e}^{2\pi{\rm i}my}$ at $y=0$, the generating function in Eq. (32) gives
$f(y)=\sum_{m=1}^{\infty}A_{m}{\rm e}^{2\pi{\rm i}my}$ $\sim\sum_{m=1}^{\infty}A_{m}\sum_{n=0}^{\infty}\frac{(2\pi{\rm i})^{n}}{n!}m^{n}y^{n}$ (40)
`````` ` ="$\sum_{n=0}^{\infty}\sum_{m=1}^{\infty}\frac{A_{m}}{m^{-n}}\frac{(2\pi{\rm i})^{n}}{n!}y^{n}$
$=\sum_{n=0}^{\infty}L(-n)\frac{(2\pi{\rm i})^{n}}{n!}y^{n}\,,$
where we have changed the order of summation. [^4] The change of summation order is a formal operation here as the coefficients $c_{n}$ might be factorially divergent. Finally, we make sense of $L(-n)$ through the analytic continuation of $L(s)$. ∎
Remark 3.1.
Depending on the explicit form of the functional equation satisfied by the $L$-function $L(s)$ in Eq. (37), it might happen that $L(-n)$ is not well-defined for some $n\in{\mathbb{Z}}_{\geq 0}$. In this case, the computation of the corresponding perturbative coefficients $c_{n}$ through the formula in Eq. (39) requires a separate ad-hoc analysis. A complete first-principles understanding of the domain of validity of Eq. (39) necessitates a specific investigation of the role played by the parameters entering the functional equation, its central point, and the motivic weight of the $L$-function. We leave this endeavour for the future.
We can now apply the propositions above to build the new modular resurgence paradigm. To do so, we assemble the key ingredients in this construction and the net of relations among them along the same lines of [20, Section 3]. Let us give a step-by-step overview below.
  • Let $\tilde{f}_{\infty}(y)\in{\mathbb{C}}[\![y]\!]$ be a Gevrey-1 formal power series whose perturbative coefficients can be written in the form of Eq. (24) with $\zeta_{m}={\cal E}_{0}m$, $m\in{\mathbb{Z}}_{\neq 0}$, for some constant ${\cal E}_{0}\in{\mathbb{C}}$, and $S_{m}\in{\mathbb{C}}$. By Proposition 3.1, $\tilde{f}_{\infty}(y)$ is a simple resurgent series whose Borel transform has a single tower of simple poles at the locations $\zeta_{m}$ with Stokes constants $S_{m}$.
  • Assume that the complex numbers $S_{m}$, $m\in{\mathbb{Z}}_{\neq 0}$, are the coefficients of two $L$-series $L_{0,+}(s)$ and $L_{0,-}(s)$, where $\{\Re(s)>\alpha\geq 0\}\subset{\mathbb{C}}$, defined as in Eq. (31). Following Proposition 3.3, we can build their generating function $f_{0}(y)$, $y\in{\mathbb{C}}\setminus{\mathbb{R}}$, defined as in Eq. (32), by taking the inverse Mellin transform of $L_{0,\pm}(s)$ as in Eq. (34).
  • Assume that $L_{0,\pm}(s)$ can be meromorphically continued to $\{\Re(s)<0\}\subset{\mathbb{C}}$. Let $\tilde{f}_{0,\pm}(y)\in{\mathbb{C}}[\![y]\!]$ be the asymptotic series of $f_{0}(y)$ as $y\to 0$ with $\Im(y)>0$ and $\Im(y)<0$, respectively. Following Proposition 3.4, the perturbative coefficients of $\tilde{f}_{0,\pm}(y)$ are explicitly given by the values of the $L$-functions $L_{0,\pm}(s)$ at negative integers as in Eq. (39), respectively. Note that $L_{0,\pm}(-n)$ for $n\in{\mathbb{Z}}_{>0}$ are computed through the functional equations that give the analytic continuation of the corresponding $L$-series. Sometimes, the analytic continuation of an $L$-function is described by another $L$-function with different coefficients. Assume this is the case and denote the new coefficients by $R_{m}\in{\mathbb{C}}$, $m\in{\mathbb{Z}}_{\neq 0}$.
  • Crucially, the resurgence structures of $\tilde{f}_{0,\pm}(y)$ are completely determined by the functional equations satisfied by $L_{0,\pm}(s)$, respectively. Assume that the perturbative coefficients of $\tilde{f}_{0}(y):=\tilde{f}_{0,+}(y)-\tilde{f}_{0,-}(y)$ can be written in terms of $\eta_{m}$ and $R_{m}$ in the form of Eq. (24), where $\eta_{m}={\cal E}_{\infty}m$, $m\in{\mathbb{Z}}_{\neq 0}$, for some constant ${\cal E}_{\infty}\in{\mathbb{C}}$. By Proposition 3.1, $\tilde{f}_{0}(y)$ is a simple resurgent series whose Borel transform has a single tower of simple poles at the locations $\eta_{m}$ with Stokes constants $R_{m}$.
  • Let $L_{\infty,\pm}(s)$, where $\{\Re(s)>\alpha\geq 0\}\subset{\mathbb{C}}$, be the two $L$-series, defined as in Eq. (31), with coefficients $R_{m}$, $m\in{\mathbb{Z}}_{\neq 0}$. Following Proposition 3.3, we can build their generating function $f_{\infty}(y)$, $y\in{\mathbb{C}}\setminus{\mathbb{R}}$, defined as in Eq. (32), by taking the inverse Mellin transform of $L_{\infty,\pm}(s)$ as in Eq. (34).
  • Finally, let $\tilde{f}_{\infty,\pm}(y)\in{\mathbb{C}}[\![y]\!]$ be the asymptotic series of $f_{\infty}(y)$ as $y\to 0$ with $\Im(y)>0$ and $\Im(y)<0$, respectively. Again following Proposition 3.4, the perturbative coefficients of $\tilde{f}_{\infty,\pm}(y)$ are explicitly given by the values of the $L$-functions $L_{\infty,\pm}(s)$ at negative integers as in Eq. (39), respectively. Once more, $L_{\infty,\pm}(-n)$ for $n\in{\mathbb{Z}}_{>0}$ are computed through the functional equations that give the analytic continuation of the corresponding $L$-series $L_{\infty,\pm}(s)$. By construction, these are determined by $L_{0,\pm}(s)$. We recover $\tilde{f}_{\infty}(y)$ as the combination $\tilde{f}_{\infty}(y)=\tilde{f}_{\infty,+}(y)-\tilde{f}_{\infty,-}(y)$.
Notice that the above relations among the dual resurgent structures of the pair of asymptotic series $\tilde{f}_{0}$, $\tilde{f}_{\infty}$ assemble into a unique global construction that generalizes the strong-weak symmetry of [20]. For instance, the $L$-functions $L_{0,\pm}(s)$ and $L_{\infty,\pm}(s)$ are are called resurgent $L$-functions in [20, Section 3]. We summarize the previous circular argument in the commutative diagram below.
[…]
Notably, starting from any fixed vertex, all other vertices of this symmetric diagram are spanned by following its directed arrows—that is, the information contained in each vertex reconstructs the whole diagram and is therefore equivalent to the information content of every other vertex. Moreover, a special place is occupied by the two-headed diagonal arrow representing the functional equation connecting the two pairs of $L$-functions $L_{0,\pm}(s)$ and $L_{\infty,\pm}(s)$.
From the conventional point of view in resurgence, looking at the generating series of the Stokes constants is unusual, as one would not generally expect them to define a function—except in those examples where there are finitely many Stokes constants, whose generating series is indeed a polynomial. However, we showed that when the resurgent structure is simple enough, and the Stokes constants have special analytic and number-theoretic properties, studying their generating series allows us to construct a new formal power series whose resurgent structure is deeply related to one of the original formal power series. In this case, the modular resurgence paradigm applies and is governed by the corresponding functional equation. Therefore, we advocate a new perspective on resurgence that aims to investigate the properties of the Stokes constants themselves.
3.2 Modular resurgent structures
Our previous discussion highlights the leading role of the $L$-functions and their functional equation in describing new features of a particular type of resurgent series. In this section, we address the relationship between quantum modularity and resurgent structures, which will clarify our choice of name for the paradigm of modular resurgence. We start by proposing the following definitions of modular resurgent structures and modular resurgent series.
Definition 3.1.
A Gevrey-1 asymptotic series $\tilde{f}(y)\in{\mathbb{C}}\llbracket y\rrbracket$ has a modular resurgent structure if the following conditions hold.
  1. 1.
    The Borel transform ${\cal B}[\tilde{f}](\zeta)\in{\mathbb{C}}\{\zeta\}$ has a tower of singularities at the locations $\zeta_{m}={\cal E}m$, $m\in\mathbb{Z}_{\neq 0}$, for some constant ${\cal E}\in{\mathbb{C}}$, in the complex $\zeta$-plane.
  2. 2.
    For every $m\in{\mathbb{Z}}_{\neq 0}$, the resurgent series at the singularity $\zeta_{m}$ is the constant function $A_{m}\in{\mathbb{C}}$, i.e., the Stokes constant.
  3. 3.
    The Stokes constants $A_{m}$, $m\in{\mathbb{Z}}_{\neq 0}$, are the coefficients of two $L$-series
    $L_{+}(s)=\sum_{m>0}\frac{A_{m}}{m^{s}}\,,\quad L_{-}(s)=-\sum_{m>0}\frac{A_{-m}}{m^{s}}$ (42)
    that are convergent in some right half-plane $\{\Re(s)>\alpha\geq 0\}\subset{\mathbb{C}}$ and can be analytically continued to $\{\Re(s)<0\}\subset{\mathbb{C}}$.
Definition 3.2.
A Gevrey-1 asymptotic series with a modular resurgent structure is a modular resurgent series. The Borel transform of a modular resurgent series is a modular resurgent function.
Note that, in the definition of modular resurgent structure, we have not specified the type of singularities appearing in the Borel plane. Indeed, their nature only affects the expression of the coefficients of the asymptotic series $\tilde{f}(y)$. For instance, if all singularities are simple poles, the perturbative coefficients depend on the non-perturbative data by means of the exact large-order relations in Eq. (24).
It follows from Proposition 3.3 that a modular resurgent series is equivalently characterized by the generating function $f(y)$ in Eq. (32). Moreover, the discontinuities across the Stokes rays identified by the tower of singularities are given by
$\mathrm{disc}_{\theta}\,\tilde{f}_{+}(y)$ $=\sum_{m=1}^{\infty}A_{m}\,{\rm e}^{-{\cal E}m/y}=f\left(-\tfrac{{\cal E}}{2\pi{\rm i}y}\right)\,,\quad y\in{\mathbb{H}}\cap\{\Re({\rm e}^{-{\rm i}\theta}y)>0\}\,,$ (43a)
$\mathrm{disc}_{2\pi-\theta}\,\tilde{f}_{-}(y)$ $=\sum_{m=1}^{\infty}A_{-m}\,{\rm e}^{{\cal E}m/y}=-f\left(-\tfrac{{\cal E}}{2\pi{\rm i}y}\right)\,,\quad y\in{\mathbb{H}}_{-}\cap\{\Re({\rm e}^{-{\rm i}\theta}y)>0\}\,,$ (43b)
where $0\leq\theta<\pi$ is the argument of the singularities in the upper half of the Borel plane, and $\tilde{f}_{\pm}(y)$ are the asymptotic series of $f(y)$ as $y\to 0$ with $\Im(y)>0$ and $\Im(y)<0$, respectively.
Let us now state our main conjectures on the modular resurgent series. Specifically, a $q$-series $f\colon{\mathbb{H}}\to{\mathbb{C}}$ whose asymptotic series at zero has a modular resurgent structure is expected to agree with the median resummation of its asymptotic expansion and to be a holomorphic quantum modular form.
Conjecture 1.
Let $f\colon{\mathbb{H}}\to{\mathbb{C}}$ be a $q$-series where $q={\rm e}^{2\pi{\rm i}y}$. If its asymptotic expansion $\tilde{f}(y)$ as $y\to 0$ with $\Im(y)>0$ has a modular resurgent structure, then the median resummation of $\tilde{f}(y)$ reconstructs the original function $f(y)$, that is,
$\mathcal{S}_{\theta}^{\mathrm{med}}\tilde{f}(y)=f(y)\,,\quad y\in{\mathbb{H}}\cap\{\Re({\rm e}^{-{\rm i}\theta}y)>0\}\,,$ (44)
where $\theta$ is the argument of the singularities in the Borel plane.
Conjecture 2.
Let $f\colon{\mathbb{H}}\to{\mathbb{C}}$ be a $q$-series where $q={\rm e}^{2\pi{\rm i}y}$. If its asymptotic expansion $\tilde{f}(y)$ as $y\to 0$ with $\Im(y)>0$ has a modular resurgent structure, then the function $f(y)$ is a holomorphic quantum modular form for a subgroup $\Gamma\subseteq\mathsf{SL}_{2}(\mathbb{Z})$.
We can relax the assumptions in the conjectures above as in Zagier’s strange identity [21]. Namely, let $f^{\prime}\colon{\mathbb{Q}}\to{\mathbb{C}}$ be such that there exists a $q$-series $f\colon{\mathbb{H}}\to{\mathbb{C}}$ whose radial limit agrees with $f^{\prime}$ to all orders. If the asymptotic expansion $\tilde{f}(y)$ as $y\to 0$ with $\Im(y)>0$ has a modular resurgent structure, then we expect that the median resummation of $\tilde{f}(y)$ reconstructs the original function $f^{\prime}(y)$, that is,
$\mathcal{S}_{\theta}^{\mathrm{med}}\tilde{f}(y)=f^{\prime}(y)\,,\quad y\in{\mathbb{Q}}\cap\{\Re({\rm e}^{-{\rm i}\theta}y)>0\}\,,$ (45)
where $\theta$ is again the argument of the singularities in the Borel plane. At the same time, we expect that $f^{\prime}(y)$ is a quantum modular form for a subgroup $\Gamma\subseteq\mathsf{SL}_{2}(\mathbb{Z})$. Moreover, note that the generating function of the Stokes constants in Eq. (32) is supported on ${\mathbb{C}}\setminus{\mathbb{R}}$. Indeed, all statements above can be recast for a $q$-series $f\colon{\mathbb{H}}_{-}\to{\mathbb{C}}$, where ${\mathbb{H}}_{-}$ denotes the lower half of the complex plane.
Furthermore, we consider functions that are defined only over the rationals. This is, in fact, the setting where quantum modular forms were originally introduced [6], as briefly discussed in Section 2.2. Conjectures 1 and 2 then become the following.
Conjecture 3.
Let $f\colon{\mathbb{Q}}\to{\mathbb{C}}$ be a $q$-series where $q={\rm e}^{2\pi{\rm i}y}$. If its asymptotic expansion $\tilde{f}(y)$ as $y\to 0$ extends to $y\in{\mathbb{C}}$ and has a modular resurgent structure, then the median resummation of $\tilde{f}(y)$ reconstructs the original function $f(y)$, that is,
$\mathcal{S}_{\theta}^{\mathrm{med}}\tilde{f}(y)=f(y)\,,\quad y\in{\mathbb{Q}}\cap\{\Re({\rm e}^{-{\rm i}\theta}y)>0\}\,,$ (46)
where $\theta$ is the argument of the singularities in the Borel plane.
Conjecture 4.
Let $f\colon{\mathbb{Q}}\to{\mathbb{C}}$ be a $q$-series where $q={\rm e}^{2\pi{\rm i}y}$. If its asymptotic expansion $\tilde{f}(y)$ as $y\to 0$ extends to $y\in{\mathbb{C}}$ and has a modular resurgent structure, then the function $f(y)$ is a quantum modular form for a subgroup $\Gamma\subseteq\mathsf{SL}_{2}(\mathbb{Z})$.
Remark 3.2.
There are (holomorphic) quantum modular forms that do not give rise to modular resurgent series. Indeed, these are the ones whose cocycle $h_{\gamma}$ has a convergent asymptotic expansion at $\gamma^{-1}(\infty)$ for all $\gamma\in\Gamma$. For instance, the cocycle of function $\tilde{\Delta}$ in Eq. (23) is a polynomial—hence, its resurgent structure is trivial.
We stress that, to perform a resurgent analysis of $\tilde{f}(y)$ in Conjectures 3 and 4, it is necessary to complexify the variable $y$. Namely, $\tilde{f}(y)$ must be analytically continued to $y\in{\mathbb{C}}$. In Conjecture 3, the median resummation $\mathcal{S}_{\theta}^{\mathrm{med}}\tilde{f}(y)$ is then restricted to the original domain of the function $f(y)$. We conclude by listing the currently available evidence in support of our conjectures.
  • In support of Conjectures 1 and 2, we have the generating functions of the weak and strong coupling Stokes constants for $\log\mathrm{Tr}(\mathsf{\rho}_{{\mathbb{P}}^{2}})$ [20] (see also Section 4.1), which are expressible in terms of $q$-Pochhammer symbols, and a collection of examples built from the theory of Maass cusp forms and their periods following Theorems 4.3 and 4.2. If we relax the assumption of $f$ being a $q$-series, then more examples are built from cusp forms in Theorem 4.4.
  • In support of Conjecture 3, we have the Kontsevich–Zagier series [^5] In the example of the Kontsevich–Zagier series, the resurgent structure has a half-tower of fractional power singularities, and not simple poles, contrary to what occurs in the examples discussed in this paper. [1] and the WRT invariants for a family of Seifert fibered homology spheres [2, Theorem 3].
  • In support of Conjecture 4, we cite again the Kontsevich–Zagier series, whose quantum modularity was proven in [21], the LMO ($\hat{Z}$) invariants for a family of plumbed 3-manifolds [13, 14], and the $q$-series $\sigma$ and $\sigma^{*}$, which are quantum modular forms according to [6, Example 1].
According to [1, Conjecture 1.4], other examples satisfying Conjecture 3 might originate from the quantum invariants of knots and 3-manifolds. For instance, the Kashaev invariants of hyperbolic knots are conjecturally vector-valued quantum modular forms [7] whose resurgent structure displays a full-fledged peacock pattern [35, 36]. Peacock patterns are also conjectured to appear in the resurgence of the fermionic spectral traces of toric CY threefolds according to [37, 19]. It would be interesting to investigate examples exhibiting multiple towers of singularities in the Borel plane and possibly extend our definition of modular resurgent structures accordingly. We plan to explore this direction in the future.
The role of $L$-functions
As we have pointed out, the modular resurgence paradigm is governed by the functional equation for the $L$-functions defined by the Stokes constants. However, at this stage, more assumptions may be needed to fully characterize these $L$-functions and enable us to prove the modular resurgence conjectures of Section 3.2. Let us collect here some considerations in this direction.
Firstly, Proposition 3.4 implies that the asymptotic series of the generating function of the Stokes constants is governed by the functional equation satisfied by the corresponding $L$-function. This is the core of the modular resurgence paradigm and explains why, starting from a given $L$-function, one might find a new divergent formal power series with a new set of Stokes constants whose resurgent structure is, however, determined by the original $L$-function. See the diagram in Eq. (41). In some examples, the functional equations for the $L$-functions $L_{\pm}(s)$ in Eq. (31) are written in terms of the same $L$-functions—as it is the case of Eq. (69). Equivalently, it might happen that the asymptotic expansion in the limit $y\to 0$ with $\Im(y)>0$ (respectively, $\Im(y)<0$) of the generating function $f(y)$ of a set of Stokes constants $\{A_{m}\}$, $m\in{\mathbb{Z}}_{\neq 0}$, which are obtained by the resurgence of the Gevrey-1 formal power series $\tilde{f}_{+}(y)$ and $\tilde{f}_{-}(y)$, is equal to $\tilde{f}_{+}(y)$ (respectively, $\tilde{f}_{-}(y)$). In this case, the commutative diagram in Eq. (41) reduces to the one below.
[…]
This simplified version of the modular resurgence construction is observed, among others, in the examples of the series $\sigma$ and $\sigma^{*}$, whose $L$-function was first studied by Cohen in [38]. [^6] The study of the resurgent structure of $\sigma$ and $\sigma^{*}$ was presented in [24, from min. 21.55]. From a number-theoretic point of view, these examples were studied in relation to quantum modular forms by Zagier [6]. More generally, a similar behavior can be observed when $f(y)$ is the function associated with a Maass cusp form of the kind considered in [39]. See Proposition 4.1.
Secondly, we might wonder which $L$-functions can appear within the modular resurgence framework. For instance, can $L$-functions coming from modular forms occur? As we will see in Section 4.3, when $L_{g}(s)$ is the $L$-function associated with a general cusp form $g:{\mathbb{H}}\to{\mathbb{C}}$, the generating function in Eq. (32) is the Fourier series of $g(y)$ itself. However, the asymptotic expansion of the cusp form for $y\to 0$ gives a convergent power series—in other words, the resurgent structure is trivial. Notably, adding to $g$ an appropriate correction that is uniquely and explicitly determined by its Fourier coefficients, we construct a function $f:{\mathbb{H}}\to{\mathbb{C}}$. Expanding the latter in the limit $y\to 0$, we obtain a Gevrey-1 asymptotic series $\tilde{f}(y)\in{\mathbb{C}}[\![y]\!]$. In this case, the simplified modular resurgence diagram in Eq. (47) does not close as shown below.
[…]
This construction tells us that $L$-functions coming from general cusp forms cannot appear in the modular resurgence paradigm. However, as we prove in Section 4.3, the above function $f$ satisfies Conjectures 1 and 2. Notice that not all modular resurgent series satisfy the modular resurgence paradigm.
4 Examples of modular resurgence
In this section, we describe two complete examples of the new paradigm of modular resurgence. In particular, we show how the global exact resurgent symmetry of Section 3.1 reproduces the strong-weak symmetry discovered in [20, Section 3] once specialized to the spectral trace of local ${\mathbb{P}}^{2}$. Then, we prove our conjectural statements in Section 3.2 for a large class of modular resurgent series originating from the theory of Maass cusp forms. Finally, in Section 4.3, we present an example of a modular resurgent structure constructed from more general cusp forms.
4.1 From topological strings
Let us compare the paradigm of modular resurgence with the strong-weak resurgent symmetry of the spectral trace of local ${\mathbb{P}}^{2}$, showing that the construction presented here and the one of [20, Section 3] are coincident and complementary.
We start by recalling the necessary background from [19, 20]. Following the prescription of [25, 40], quantizing the mirror curve to local ${\mathbb{P}}^{2}$ gives rise to a quantum operator acting on $L^{2}({\mathbb{R}})$ whose inverse $\rho_{{\mathbb{P}}^{2}}$ is positive-definite and of trace class. Its first fermionic spectral trace [^7] The fermionic spectral traces $Z_{X}(\bm{N},\hbar)$ provide a non-perturbative definition of the topological string on a toric CY threefold $X$ according to the conjectural Topological String/Spectral Theory (TS/ST) correspondence. The Planck constant $\hbar$ satisfies a strong-weak coupling duality $\hbar\propto g_{s}^{-1}$, where $g_{s}$ is the string coupling constant. $Z_{{\mathbb{P}}^{2}}(1,\hbar)=\mathrm{Tr}(\mathsf{\rho}_{{\mathbb{P}}^{2}})$ is a well-defined analytic function of the quantum deformation parameter $\hbar\in{\mathbb{C}}^{\prime}$, which is known in closed form for $\Im(\hbar)>0$ [41]. The all-orders perturbative expansions of the logarithm of the spectral trace in the limits of $\hbar\rightarrow 0$ and $\hbar\rightarrow\infty$ have been computed by the second author in [19]. They give rise to the Gevrey-1 asymptotic series
ϕ()italic-$\phi(\hbar)$ $=\sum_{n=1}^{\infty}a_{2n}\hbar^{2n}\in{\mathbb{Q}}[\![\hbar]\!]\,,$ (49a)
$\psi(\tau)$ $=\sum_{n=1}^{\infty}b_{2n}\tau^{2n-1}\in{\mathbb{Q}}[\pi,\sqrt{3}][\![\tau]\!]\,,$ (49b)
where $\tau=-\tfrac{2\pi}{3\hbar}$, respectively. Their Borel transforms possess a single tower of simple poles at the locations
$\zeta_{m}={\cal A}_{0}{\rm i}m\,,\quad\eta_{m}={\cal A}_{\infty}{\rm i}m\,,\quad m\in{\mathbb{Z}}_{\neq 0}\,,$ (50)
where ${\cal A}_{0}=\frac{4\pi^{2}}{3}$ and ${\cal A}_{\infty}=\frac{2\pi}{3}$, respectively. The corresponding Stokes constants
$S_{m}=S_{-m}\in\sqrt{3}{\rm i}\,{\mathbb{Q}}_{>0}\,,\quad R_{m}=-R_{-m}\in{\mathbb{Q}}_{\neq 0}\,,\quad m\in{\mathbb{Z}}_{\neq 0}\,,$ (51)
are known explicitly as divisor sum functions and define the weak and strong coupling $L$-series
$L_{0}(s)=\sum_{m=1}^{\infty}\frac{S_{m}}{m^{s}}\,,\quad L_{\infty}(s)=\sum_{m=1}^{\infty}\frac{R_{m}}{m^{s}}\,,$ (52)
which converge in the right half-plane $\Re(s)>1$ and can be analytically continued to meromorphic functions on the whole complex $s$-plane [19, Section 4.4]. Therefore, ϕ()italic-$\phi(\hbar)$ and $\psi(\tau)$ are modular resurgent series according to Definition 3.1.
In the companion paper [20], we complete the resurgent $L$-functions in Eq. (52) into
$\Lambda_{0}(s)$ $=\frac{3^{\frac{s}{2}-1}}{{\rm i}\pi^{s+1}}\Gamma\left(\frac{s}{2}\right)\Gamma\left(\frac{s}{2}+1\right)L_{0}(s)\,,$ (53a)
$\Lambda_{\infty}(s)$ $=\frac{3^{\frac{s}{2}-1}}{\pi^{s+1}}\Gamma\left(\frac{s+1}{2}\right)^{2}L_{\infty}(s)\,,$ (53b)
and show that $\Lambda_{0}(s)$, $\Lambda_{\infty}(s)$ are analytically continued to the whole complex $s$-plane through each other. In particular, they satisfy the functional equation
$\Lambda_{0}(s)=\Lambda_{\infty}(-s)\,.$ (54)
Simultaneously, we take the perspective of the generating functions $f_{0}(y)$, $f_{\infty}(y)$, $y\in{\mathbb{C}}\setminus{\mathbb{R}}$, of the weak and strong coupling Stokes constants, which satisfy the parity properties
$f_{0}(-y)=-f_{0}(y)\,,\quad f_{\infty}(-y)=f_{\infty}(y)\,.$ (55)
Their all-orders perturbative expansions in the limit $y\rightarrow 0$ with $\Im(y)>0$ are
$\tilde{f}_{0}(y)$ $=-\frac{\pi{\rm i}}{2}-\frac{3\mathcal{V}}{2\pi{\rm i}y}-2\psi(y)\,,$ (56a)
$\tilde{f}_{\infty}(y)$ $=-3\log\frac{\Gamma(2/3)}{\Gamma(1/3)}-\log(-6\pi{\rm i}y)+2\phi(2\pi y)\,,$ (56b)
where $\mathcal{V}=2\Im\left(\mathrm{Li}_{2}({\rm e}^{2\pi{\rm i}/3})\right)$.
Let us neglect the terms that do not enter the perturbative series in the RHS of Eqs. (56a) and (56b). We show that they satisfy the newly introduced paradigm of modular resurgence. This offers a complementary understanding of the strong-weak resurgent symmetry discussed in [20, Section 3], particularly highlighting the central role played by the functional equation in Eq. (54). For completeness, we repeat the main steps in the modular resurgence construction for this key example.
  • The resurgent structure of the asymptotic series $\tilde{f}_{\infty}(y)$ is dictated by the formal power series $2\phi(2\pi y)$ according to the formula in Eq. (56b). Thus, the resurgent structure is modular and can be written in the coordinate conjugate to $y$ as $\{\eta_{m},\,S_{m}\}$, $m\in{\mathbb{Z}}_{\neq 0}$, where the Stokes constants satisfy $S_{-m}=S_{m}$.
  • The weak coupling $L$-series $L_{0}(s)$, $s\in{\mathbb{C}}$ with $\Re(s)>1$, has coefficients given by the Stokes constants $S_{m}$ for $m\in{\mathbb{Z}}_{>0}$. Note that $L_{0}(s)=\pm L_{0,\pm}(s)$ in the notation of Eq. (31). Because of the parity of the Stokes constants, we need only consider one $L$-series.
  • Taking the inverse Mellin transform of $L_{0}(s)$ as in Eq. (34), we obtain the generating function $f_{0}(y)$. Let
    $\tilde{f}_{0}(y)=\sum_{n=1}^{\infty}c_{n}^{0}y^{n}\in{\mathbb{C}}\llbracket y\rrbracket$ (57)
    be its asymptotic series as $y\to 0$ with $\Im(y)>0$. By Proposition 3.4, particularly Eq. (39), the perturbative coefficients $c_{n}^{0}$, $n\in{\mathbb{Z}}_{>0}$, are given by
    $c_{n}^{0}=L_{0}(-n)\frac{(2\pi{\rm i})^{n}}{n!}\,,$ (58)
    where we make sense of the RHS via the functional equation for the completed $L$-function $\Lambda_{0}(s)$. In particular, substituting Eqs. (53a) and (53b) into Eq. (54), we find that
    $L_{0}(-s)=-\frac{2{\rm i}3^{s}}{\pi^{2s}s}\frac{\Gamma\left(\frac{s+1}{2}\right)^{2}}{\Gamma\left(-\frac{s}{2}\right)^{2}}L_{\infty}(s)\,.$ (59)
    Notice that $L_{0}(-s)$ tends to zero for $s=n\in{\mathbb{Z}}_{>0}$ even as a consequence of the divergence of the gamma function in the denominator. Thus, we set $c_{2n}^{0}=0$, $n\in{\mathbb{Z}}_{>0}$. Then, taking $s=n\in{\mathbb{Z}}_{>0}$ odd, applying the well-known properties of the gamma function, and using the series representation of $L_{\infty}(s)$ in Eq. (52), Eq. (59) becomes
    $L_{0}(-n)$ $=-\frac{2{\rm i}}{\pi}n!(n-1)!(-2\pi{\rm i})^{-n}\sum_{m=1}^{\infty}\frac{R_{m}}{\eta_{m}^{n}}\,.$ (60)
    Applying Eq. (58), the non-trivial perturbative coefficients $c_{2n+1}^{0}$, $n\in{\mathbb{Z}}_{\geq 0}$, are
    $c_{2n+1}^{0}=-2\frac{\Gamma(2n+1)}{\pi{\rm i}}\sum_{m=1}^{\infty}\frac{R_{m}}{\eta_{m}^{2n+1}}=-2b_{2n+2}\,,$ (61)
    where we have used the exact large-order identity in Eq. (24) for the perturbative coefficients $\{b_{2k}\}$, $k\in{\mathbb{Z}}_{>0}$, of the formal power series $\psi(\tau)$ in Eq. (49b). This is what we expect from the formula in Eq. (56a).
  • The resurgent structure of the asymptotic series $\tilde{f}_{0}(y)$ is dictated by the formal power series $-2\psi(y)$ according to the formula in Eq. (56a). Thus, the resurgent structure is modular and can be written in the coordinate conjugate to $y$ as $\{\eta_{m},R_{m}\}$, $m\in{\mathbb{Z}}_{\neq 0}$, where the Stokes constants satisfy $R_{-m}=-R_{m}$.
  • The strong coupling $L$-series $L_{\infty}(s)$, $s\in{\mathbb{C}}$ with $\Re(s)>1$, has coefficients given by the Stokes constants $R_{m}$ for $m\in{\mathbb{Z}}_{>0}$. Note that $L_{\infty}(s)=L_{\infty,\pm}(s)$ in the notation of Eq. (31). Again, because of the parity of the Stokes constants, we need only consider one $L$-series.
  • Taking the inverse Mellin transform of $L_{\infty}(s)$ as in Eq. (34), we obtain the generating function $f_{\infty}(y)$. Let
    $\tilde{f}_{\infty}(y)=\sum_{n=1}^{\infty}c_{n}^{\infty}y^{n}\in{\mathbb{C}}\llbracket y\rrbracket$ (62)
    be its asymptotic series as $y\to 0$ with $\Im(y)>0$. By Proposition 3.4, particularly Eq. (39), its perturbative coefficients $c_{n}^{\infty}$, $n\in{\mathbb{Z}}_{>0}$, are given by
    $c_{n}^{\infty}=L_{\infty}(-n)\frac{(2\pi{\rm i})^{n}}{n!}\,,$ (63)
    where we make sense of the RHS via the functional equation for the completed $L$-function $\Lambda_{\infty}(s)$ as before. In particular, Eq. (59) gives
    $L_{\infty}(-s)=\frac{3^{s}s}{2{\rm i}\pi^{2s}}\frac{\Gamma\left(\frac{s}{2}\right)^{2}}{\Gamma\left(\frac{1-s}{2}\right)^{2}}L_{0}(s)\,.$ (64)
    Notice that $L_{\infty}(-s)$ tends to zero for $s=n\in{\mathbb{Z}}_{\geq 0}$ odd due to the divergence of the gamma function in the denominator. Hence, we set $c_{2n+1}^{\infty}=0$, $n\in{\mathbb{Z}}_{\geq 0}$. Then, taking $s=n\in{\mathbb{Z}}_{>0}$ even, applying the well-known properties of the gamma function, and using the series representation of $L_{0}(s)$ in Eq. (52), Eq. (64) becomes
    $L_{\infty}(-n)$ $=\frac{2}{{\rm i}\pi}n!(n-1)!(-2\pi)^{n}\sum_{m=1}^{\infty}\frac{S_{m}}{\zeta_{m}^{n}}\,.$ (65)
    Applying Eq. (63), the non-trivial perturbative coefficients $c_{2n}^{\infty}$, $n\in{\mathbb{Z}}_{>0}$, are
    $c_{2n}^{\infty}=2(2\pi)^{2n}\frac{\Gamma(2n)}{\pi{\rm i}}\sum_{m=1}^{\infty}\frac{S_{m}}{\zeta_{m}^{2n}}=2(2\pi)^{2n}a_{2n}\,,$ (66)
    where we have used the exact large-order identity in Eq. (24) for the perturbative coefficients $\{a_{2k}\}$, $k\in{\mathbb{Z}}_{>0}$, of the formal power series ϕ()italic-$\phi(\hbar)$ in Eq. (49a). Note again that this is what we expect from the formula in Eq. (56b).
Remark 4.1.
The term of order $y^{0}$ does not enter our definitions of $\tilde{f}_{0}(y)$ and $\tilde{f}_{\infty}(y)$. We observe, however, that direct evaluation of the weak coupling $L$-function $L_{0}(s)$ at $s=0$ using the closed formula in [19, Eq. (4.174a)] gives the perturbative coefficient $c_{0}^{0}=L_{0}(0)=-\pi{\rm i}/2$ according to the formula in Eq. (58), which agrees with the asymptotic expansion in Eq. (56a). At the same time, the strong coupling $L$-function $L_{\infty}(s)$ in [19, Eq. (4.174b)] is not well-defined at $s=0$—as it can also be observed from the functional equation in Eq. (64). In particular, following Eq. (63), the perturbative coefficient $c_{0}^{\infty}$ would appear to diverge to infinity. We conclude that the constant term in the asymptotic expansion in Eq. (56b) cannot be consistently derived from the analytic continuation of the $L$-function $L_{\infty}(s)$.
We have shown that the weak and strong coupling resurgent structures of the spectral trace of local ${\mathbb{P}}^{2}$ are a modular resurgent pair and satisfy the global symmetry of Section 3.1. Remarkably, as we argue in [20], this construction is responsible for the exchange of the perturbative/non-perturbative contributions to the holomorphic and anti-holomorphic blocks in the well-known factorization of the spectral trace, which can be intuitively traced back to the duality between the worldsheet and WKB contributions to the total grand potential of the topological string on local ${\mathbb{P}}^{2}$.
Finally, we stress that both Conjectures 1 and 2 are explicitly proven for the generating function $f_{0}$. Namely, $f_{0}$ is a holomorphic quantum modular function for the congruence subgroup $\Gamma_{1}(3)\subset\mathsf{SL}_{2}({\mathbb{Z}})$ [20, Theorem 4.6] and is reconstructed by the median resummation of its asymptotic expansion [20, Theorem 4.7]. As for the generating function $f_{\infty}$, Conjecture 2 is proven with respect to the same group $\Gamma_{1}(3)$ [20, Theorem 4.7], while Conjecture 1 is only supported by numerical evidence [20, Conjecture 1].
4.2 From Maass cusp forms
Given a pair of $L$-series satisfying special functional equations, a large class of examples of modular resurgent series can be derived from this data. In particular, building on the results of [39], we consider $L$-functions related to Maass cusp forms with spectral parameter 1/2121/21 / 2.888A Maass cusp form with spectral parameter $s\in{\mathbb{C}}$ is a smooth $\mathsf{PSL}_{2}({\mathbb{Z}})$-invariant function $u\colon{\mathbb{H}}\to{\mathbb{C}}$ that is an eigenfunction of the hyperbolic Laplacian $\Delta$ with eigenvalue $s(1-s)$. Namely, $\Delta u=s(1-s)u$. These functions give a basis for $L^{2}$ on ${\mathbb{H}}/\mathsf{PSL}_{2}({\mathbb{Z}})$.
Proposition 4.1.
Let $\{A_{m}\}$, $m\in{\mathbb{Z}}_{\neq 0}$, be a sequence of complex numbers such that the two $L$-series
$L_{\epsilon}(s)=\sum_{m=1}^{\infty}\frac{A_{m}+(-1)^{\epsilon}A_{-m}}{m^{s}}\,,\quad\epsilon=0,1\,,$ (67)
are convergent in some right half-plane $\{\Re(s)>\alpha\geq 0\}\subset{\mathbb{C}}$. Assume that the corresponding completed $L$-functions
$\Lambda_{\epsilon}(s)=\gamma(s+\epsilon)L_{\epsilon}(s)\,,\quad\text{where}\quad\gamma(s)=\frac{1}{4\pi^{s}}\Gamma\left(\frac{s}{2}\right)^{2}\,,$ (68)
are entire functions of finite order and satisfy the functional equations
$\Lambda_{\epsilon}(1-s)=(-1)^{\epsilon}\Lambda_{\epsilon}(s)\,,\quad\epsilon=0,1\,.$ (69)
Let $f\colon{\mathbb{C}}\setminus{\mathbb{R}}\to{\mathbb{C}}$ be the generating function
$f(y):=\begin{cases}\displaystyle\sum_{m>0}A_{m}{\rm e}^{2\pi{\rm i}my}&\text{ if }\Im(y)>0\\ &\\ -\displaystyle\sum_{m<0}A_{m}{\rm e}^{2\pi{\rm i}my}&\text{ if }\Im(y)<0\end{cases}\,.$ (70)
The asymptotic series $\tilde{f}_{+}(y)$ and $\tilde{f}_{-}(y)$ of $f(y)$ as $y\to 0$ with $\Im(y)>0$ and $\Im(y)<0$, respectively, are modular resurgent series.
Proof.
Let us introduce the two $L$-functions
$L_{+}(s)$ $=\sum_{m=1}^{\infty}\frac{A_{m}}{m^{s}}=\frac{L_{0}(s)+L_{1}(s)}{2}\,,$ (71a)
$L_{-}(s)$ $=-\sum_{m=1}^{\infty}\frac{A_{-m}}{m^{s}}=\frac{L_{1}(s)-L_{0}(s)}{2}\,,$ (71b)
as in Eq. (31). It follows from Proposition 3.4, particularly from Eq. (39), that the asymptotic series of $f(y)$ as $y\to 0$ with $\Im(y)>0$ can be written as
$\tilde{f}_{+}(y)=\sum_{n=0}^{\infty}L_{+}(-n)\frac{(2\pi{\rm i})^{n}}{n!}y^{n}\,.$ (72)
Similarly, its asymptotic series as $y\to 0$ with $\Im(y)<0$ is
$\tilde{f}_{-}(y)=\sum_{n=0}^{\infty}L_{-}(-n)\frac{(2\pi{\rm i})^{n}}{n!}y^{n}\,.$ (73)
We can now apply the functional equation for $\Lambda_{\epsilon}(s)$ in Eq. (69) to make sense of $L_{\pm}(-n)$ for $n\in{\mathbb{Z}}_{\geq 0}$. To do so, let us consider separately the case of $n\in{\mathbb{Z}}_{\geq 0}$ even (respectively, odd) for $L_{1}(-n)$ (respectively, $L_{0}(-n)$) due to the expression for the gamma factor $\gamma(s)$. For $n\in{\mathbb{Z}}_{\geq 0}$ even, we have that
$L_{1}(-n)$ $=-\frac{\gamma(n+2)}{\gamma(1-n)}L_{1}(n+1)=-\frac{2^{-2n}}{\pi^{2n+2}}\Gamma(n+1)^{2}L_{1}(n+1)\,,$ (74)
where we have used the definition for the gamma factor in Eq. (68) and applied the well-known identities for the gamma function. Similarly, for $n\in{\mathbb{Z}}_{>0}$ odd, we have that
$L_{0}(-n)$ $=\frac{\gamma(n+1)}{\gamma(-n)}L_{0}(n+1)=\frac{2^{-2n}}{\pi^{2n+2}}\Gamma(n+1)^{2}L_{0}(n+1)\,.$ (75)
Since the gamma function has simple poles at the non-positive integers, we set $L_{1}(-n)=0$ for $n\in{\mathbb{Z}}_{>0}$ odd and $L_{0}(-n)=0$ for $n\in{\mathbb{Z}}_{\geq 0}$ even. Therefore, putting together Eqs. (71a), (74), and (75), we find that
$\dfrac{1}{n!}L_{+}(-n)$ =$=\begin{cases}-\dfrac{2^{-2n-1}}{\pi^{2n+2}}n!\,L_{1}(n+1)=-\dfrac{2^{-2n-1}}{\pi^{2n+2}}n!\,\displaystyle\sum_{m=1}^{\infty}\dfrac{A_{m}-A_{-m}}{m^{n+1}}&\text{ if }n\in{\mathbb{Z}}_{\geq 0}\text{ even}\\ &\\ \dfrac{2^{-2n-1}}{\pi^{2n+2}}n!\,L_{0}(n+1)=\dfrac{2^{-2n-1}}{\pi^{2n+2}}n!\,\displaystyle\sum_{m=1}^{\infty}\dfrac{A_{m}+A_{-m}}{m^{n+1}}&\text{ if }n\in{\mathbb{Z}}_{>0}\text{ odd}\end{cases}$ (76)
$=\frac{2^{-2n-1}}{\pi^{2n+2}}n!\left(\sum_{m=1}^{\infty}\frac{A_{m}}{(-m)^{n+1}}+\sum_{m=1}^{\infty}\frac{A_{-m}}{m^{n+1}}\right)$
$=\frac{2^{-2n-1}}{\pi^{2n+2}}n!\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{(-m)^{n+1}}\,.$
Analogously, putting together Eqs. (71b), (74), and (75), we find that
$\dfrac{1}{n!}L_{-}(-n)$ =$=\begin{cases}-\dfrac{2^{-2n-1}}{\pi^{2n+2}}n!\,L_{1}(n+1)=-\dfrac{2^{-2n-1}}{\pi^{2n+2}}n!\,\displaystyle\sum_{m=1}^{\infty}\dfrac{A_{m}-A_{-m}}{m^{n+1}}&\text{ if }n\in{\mathbb{Z}}_{\geq 0}\text{ even}\\ &\\ -\dfrac{2^{-2n-1}}{\pi^{2n+2}}n!\,L_{0}(n+1)=-\dfrac{2^{-2n-1}}{\pi^{2n+2}}n!\,\displaystyle\sum_{m=1}^{\infty}\dfrac{A_{m}+A_{-m}}{m^{n+1}}&\text{ if }n\in{\mathbb{Z}}_{>0}\text{ odd}\end{cases}$ (77)
$=-\frac{2^{-2n-1}}{\pi^{2n+2}}n!\left(\sum_{m=1}^{\infty}\frac{A_{m}}{m^{n+1}}+\sum_{m=1}^{\infty}\frac{A_{-m}}{(-m)^{n+1}}\right)$
$=-\frac{2^{-2n-1}}{\pi^{2n+2}}n!\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{m^{n+1}}\,.$
To sum up, substituting Eq. (76) into Eq. (72) yields
$\tilde{f}_{+}(y)$ $=2\sum_{n=0}^{\infty}n!\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{(-4\pi^{2}m)^{n+1}}(2\pi{\rm i})^{n}y^{n}\,$ (78)
$=\sum_{n=0}^{\infty}\left(\frac{n!}{\pi{\rm i}}\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{(2\pi{\rm i}m)^{n+1}}\right)y^{n}\,,$
which is a Gevrey-1 asymptotic series. Using Proposition 3.1, we conclude that the Borel transform of $y\tilde{f}_{+}(y)$ has a modular resurgent structure with simple poles at $\rho_{m}=2\pi{\rm i}m$, $m\in{\mathbb{Z}}_{\neq 0}$, and Stokes constants $2A_{m}$. Analogously, substituting Eq. (77) into Eq. (73) gives
$\tilde{f}_{-}(y)$ $=-2\sum_{n=0}^{\infty}n!\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{(4\pi^{2}m)^{n+1}}(2\pi{\rm i})^{n}y^{n}\,$ (79)
$=-\sum_{n=0}^{\infty}\left(\frac{n!}{\pi{\rm i}}\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{-m}}{(2\pi{\rm i}m)^{n+1}}\right)y^{n}\,,$
which is a Gevrey-1 asymptotic series. Once more, using Proposition 3.1, we conclude that the Borel transform of $y\tilde{f}_{-}(y)$ has a modular resurgent structure with simple poles at $\rho_{m}=2\pi{\rm i}m$, $m\in{\mathbb{Z}}_{\neq 0}$, and Stokes constants $-2A_{-m}$. ∎
Notice that the modular resurgent series of Proposition 4.1 satisfy the simplified diagram in Eq. (47). Additionally, the generating function defined in Eq. (70) satisfies Conjecture 2.
Theorem 4.2.
Within the assumptions of Proposition 4.1, the generating function $f:{\mathbb{C}}\setminus{\mathbb{R}}\to{\mathbb{C}}$, defined in Eq. (70), restricted to either the upper or the lower half of the complex plane is a weight-$1$ holomorphic quantum modular form for the group $\mathsf{SL}_{2}({\mathbb{Z}})$.
Proof.
Recall from part $(d)$ of Theorem $1$ in [39, Chapter 1] that there exists an analytic function $\psi\colon{\mathbb{C}}^{\prime}\to{\mathbb{C}}$ such that [^9] The function $\psi$ in Eq. (80) is the so-called period of the function $f$ in Eq. (70) and satisfies a three-terms functional equation. See [39] for details.
$c\,\psi(y)=f(y)-\frac{1}{y}f\left(-\frac{1}{y}\right)\,,$ (80)
for some constant $c\in{\mathbb{C}}_{\neq 0}$. Moreover, recall that the generators of the modular group $\mathsf{SL}_{2}({\mathbb{Z}})$ are the matrices
$T=\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\,,\quad S=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}\,.$ (81)
Hence, if we restrict the generating function to the upper half-plane, i.e., we take $f\colon{\mathbb{H}}\to{\mathbb{C}}$, the cocycle $h_{T}[f](y)$ is trivial due to the periodicity property $f(y+1)=f(y)$, while the cocycle for $S$ is given by
$h_{S}[f](y)=-c\,\psi(y)\,.$ (82)
The same arguments apply to the restriction of the generating function to the lower half-plane, i.e., for $f\colon{\mathbb{H}}_{-}\to{\mathbb{C}}$. ∎
With an extra assumption on the parity of the coefficients $A_{m}$, $m\in{\mathbb{Z}}_{\neq 0}$, appearing in the definition of the $L$-functions $L_{\epsilon}(s)$, $\epsilon=0,1$, in Eq. (67), we prove the effectiveness of the median resummation for the generating function defined in Eq. (70), which therefore satisfies Conjecture 1. In particular, assuming $A_{m}=A_{-m}$, we have that $f(-y)=-f(y)$ for $y\in{\mathbb{C}}\setminus{\mathbb{R}}$, so that $f(y)$ need only be specified in the upper half of the complex $y$-plane.
Theorem 4.3.
Within the assumptions of Proposition 4.1, if $A_{m}=A_{-m}$ for every $m\in{\mathbb{Z}}_{\neq 0}$, the generating function $f\colon{\mathbb{H}}\to{\mathbb{C}}$ defined in Eq. (70) is recovered from its asymptotic expansion $\tilde{f}_{+}(y)$ as $y\to 0$ with $\Im(y)>0$ through the median resummation. Precisely,
$yf(y)=\mathcal{S}^{\mathrm{med}}_{\frac{\pi}{2}}\big{[}y\tilde{f}_{+}(y)\big{]}\,,\quad y\in{\mathbb{H}}\,.$ (83)
Proof.
Due to the parity of the coefficients, $L_{1}(s)$ is identically zero, while
$L_{0}(s)=2\sum_{m=1}^{\infty}\frac{A_{m}}{m^{s}}\,,$ (84)
so that the two $L$-functions in Eqs. (71a) and (71b) reduce to
$L_{\pm}(s)=\frac{L_{0}(s)}{2}\,.$ (85)
As a consequence of Proposition 4.1, and particularly Eqs. (78) and (79), the formal power series $\tilde{f}_{\pm}(y)$ can be written as
$\tilde{f}_{\pm}(y)=\pm\frac{2}{\pi{\rm i}}\sum_{n=0}^{\infty}(2n+1)!\sum_{m>0}\frac{A_{m}}{\rho_{m}^{2n+2}}y^{2n+1}\,,$ (86)
where $\rho_{m}=2\pi{\rm i}m$ as before. Recall that the median resummation at angle $\pi/2$ can be expressed in terms of the Borel–Laplace sums at angles 00 and $\pi$ using Eq. (15). We have that
$\mathcal{S}^{\mathrm{med}}_{\frac{\pi}{2}}\big{[}y\tilde{f}_{+}(y)\big{]}$ =$=\begin{cases}s_{0}\big{(}y\tilde{f}_{+}(y)\big{)}+\frac{1}{2}\mathrm{disc}_{\frac{\pi}{2}}\big{[}y\tilde{f}_{+}(y)\big{]}\,,&\Re(y)>0\,,\\ \\ s_{\pi}\big{(}y\tilde{f}_{+}(y)\big{)}-\frac{1}{2}\mathrm{disc}_{\frac{\pi}{2}}\big{[}y\tilde{f}_{+}(y)\big{]}\,,&\Re(y)<0\,,\end{cases}$ (87)
for $\Im(y)>0$. To compute the Borel–Laplace sums of $\tilde{f}_{+}(y)$, let us first evaluate its Borel transform as in the proof of Proposition 3.1, that is,
$\mathcal{B}\big{[}y\tilde{f}_{+}(y)\big{]}(\zeta)=\frac{2}{\pi{\rm i}}\sum_{n=0}^{\infty}\sum_{m>0}\frac{A_{m}}{\rho_{m}^{2n+2}}\zeta^{2n+1}=-\frac{1}{\pi{\rm i}}\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{\zeta-\rho_{m}}\,,$ (88)
where we have exchanged the order of summation, resummed the geometric series over the index $n$, and applied the identity
$\frac{1}{(x-y)(x+y)}=\frac{1}{2x(x-y)}+\frac{1}{2x(x+y)}$ (89)
in the second step. Then, the Borel–Laplace sums at angles 00 and $\pi$ are
$s_{0}\big{(}y\tilde{f}_{+}(y)\big{)}$ $=-\frac{1}{\pi{\rm i}}\int_{0}^{\infty}{\rm e}^{-\zeta/y}\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{\zeta-\rho_{m}}\,d\zeta=-2\sum_{m\in{\mathbb{Z}}_{\neq 0}}A_{m}\mathbf{e}_{1}\left(-\frac{m}{y}\right)\,,\quad$ $y>0\,,$ (90a)
$s_{\pi}\big{(}y\tilde{f}_{+}(y)\big{)}$ $=-\frac{1}{\pi{\rm i}}\int_{0}^{-\infty}{\rm e}^{-\zeta/y}\sum_{m\in{\mathbb{Z}}_{\neq 0}}\frac{A_{m}}{\zeta-\rho_{m}}\,d\zeta=-2\sum_{m\in{\mathbb{Z}}_{\neq 0}}A_{m}\mathbf{e}_{1}\left(-\frac{m}{y}\right)\,,\quad$ $y<0\,,$ (90b)
where we have introduced the analytic function $\mathbf{e}_{1}\colon{\mathbb{C}}\setminus{\rm i}{\mathbb{R}}_{\geq 0}\to{\mathbb{C}}$ defined by [^10] The notation in Eq. (91) was suggested by M. Kontsevich as the function $\mathbf{e}_{1}$ plays a similar role in other examples (e.g., the case of $\sigma,\sigma^{*}$ in [24, from min. 21.55] and the spectral trace of local ${\mathbb{P}}^{2}$ in [20]).
$\mathbf{e}_{1}(z):=\frac{1}{2\pi{\rm i}}\int_{0}^{\infty}{\rm e}^{-2\pi t}\frac{dt}{t+{\rm i}z}=\frac{1}{2\pi{\rm i}}{\rm e}^{2\pi{\rm i}z}\,\Gamma(0,2\pi{\rm i}z)\,,$ (91)
where $\Gamma(s,2\pi{\rm i}z)$ denotes the upper incomplete gamma function. Note that the Borel–Laplace sums above can be analytically continued to $\Re(y)>0$ and $\Re(y)<0$, respectively, as a consequence of the properties of the function $\mathbf{e}_{1}$. Besides, by the standard residue argument, the discontinuity is
$\mathrm{disc}_{\frac{\pi}{2}}\big{[}y\tilde{f}_{+}(y)\big{]}=2\sum_{m>0}A_{m}{\rm e}^{-2\pi{\rm i}m/y}=2f\left(-\frac{1}{y}\right)\,,\quad\Im(y)>0\,.$ (92)
We can then apply the results of Section 3, Chapter II of [39]. In particular, adopting the notation of [39, Eqs. (2.11) and (2.13)], we define the function $\psi_{1}\colon{\mathbb{C}}\setminus{\rm i}{\mathbb{R}}_{\geq 0}\to{\mathbb{C}}$ as
$\psi_{1}(y):=2\pi{\rm i}\sum_{m\in{\mathbb{Z}}_{\neq 0}}A_{m}\mathbf{e}_{1}(my)\,,$ (93)
which satisfies the properties [39, Eq. (2.10)]
$y\psi_{1}(y)=\psi_{1}\left(\frac{1}{y}\right)\,,\quad\psi_{1}(-y)=\psi_{1}(y)\,,$ (94)
where the first formula holds for $\Re(y)>0$. Then, we have that [39, Eq. (2.20)] [^11] The same formula in Eq. (95) holds for $y\in{\mathbb{H}}_{-}$ with a change of sign in the constant $c_{*}$.
$c_{*}\,f(y)=\psi_{1}(y)+\frac{1}{y}\psi_{1}\left(-\frac{1}{y}\right)\,,$ (95)
for some constant $c_{*}\in{\mathbb{C}}_{\neq 0}$. Let us set the constant $c_{*}=-2\pi{\rm i}$. Using Eqs. (95) and (94), we deduce that
$yf(y)$ =$=\begin{cases}\displaystyle-\frac{1}{\pi{\rm i}}\psi_{1}\left(-\frac{1}{y}\right)+f\left(-\frac{1}{y}\right)\,,\quad&\Re(y)>0\,,\\ &\\ \displaystyle-\frac{1}{\pi{\rm i}}\psi_{1}\left(-\frac{1}{y}\right)-f\left(-\frac{1}{y}\right)\,,\quad&\Re(y)<0\,.\end{cases}$ (96)
It follows from Eqs. (93), (90a), and (90b) that we can identify $\psi_{1}(-1/y)$ with the Borel–Laplace sum of $y\tilde{f}_{+}(y)$ along the positive real axis for $\Re(y)>0$, that is,
$-\frac{1}{\pi{\rm i}}\psi_{1}\left(-\frac{1}{y}\right)=s_{0}\big{(}y\tilde{f}_{+}(y)\big{)}\,,\quad\Re(y)>0\,,$ (97)
and with its Borel–Laplace sum along the negative real axis for $\Re(y)<0$, that is,
$-\frac{1}{\pi{\rm i}}\psi_{1}\left(-\frac{1}{y}\right)=s_{\pi}\big{(}y\tilde{f}_{+}(y)\big{)}\,,\quad\Re(y)<0\,.$ (98)
Substituting the discontinuity in Eq. (92) and the Borel–Laplace sums in Eqs. (97) and (98) into the formula for the median resummation in Eq. (87), we find the same expression in Eq. (96). ∎
A result analogous to Theorem 4.3 applies for an odd sequence of complex numbers $A_{m}$, $m\in{\mathbb{Z}}_{\neq 0}$, i.e., such that $A_{m}=-A_{-m}$. In this case, the generating function in Eq. (70) satisfies $f(-y)=f(y)$ for $y\in{\mathbb{C}}\setminus{\mathbb{R}}$, and $L_{1}(s)=2\sum_{m=1}^{\infty}\tfrac{A_{m}}{m^{s}}$ is the only non-trivial $L$-function. Moreover, we expect a similar result for Maass cusps forms that are invariant under a discrete subgroup of $\mathsf{PSL}_{2}({\mathbb{Z}})$.
Remark 4.2.
The example of even Maass cusp forms discussed here displays an interesting feature. Indeed, the cocycle $\psi(y)$ in Eq. (80) is proportional to the function $\psi_{1}(y)$ in Eq. (93) by means of Eq. (95). Consequently, it follows from Theorems 4.2 and 4.3 that the Borel–Laplace sums $s_{0}\big{(}y\tilde{f}_{\pm}(y)\big{)}$ are proportional to $\psi(-1/y)$. We expect this property to hold more generally when both Conjectures 1 and 2 (or, equivalently, Conjectures 3 and 4) are verified—potentially, under some additional assumptions on the $q$-series. In other words, we expect the cocycles of (a well-defined class of) $q$-series whose asymptotic expansions have a modular resurgent structure to occur as Borel–Laplace sums.
4.3 From cusp forms
In this section, we consider the $L$-function of a general cusp form. Breaking the modular invariance of the cusp form by adding to it a function written in terms of $\mathbf{e}_{1}$, which we have defined in Eq. (91), we build a holomorphic function whose asymptotics is a modular resurgent series. In particular, its Stokes constants are the Fourier coefficients of the original cusp form. In addition, we show that the given holomorphic function, although not anymore a $q$-series, gives evidence of Conjecture 1 and, after a suitable change of coordinate, of Conjecture 2.
Theorem 4.4.
Let $g\colon{\mathbb{H}}\to{\mathbb{C}}$ be a weight-$\omega$ cusp form for $\mathsf{SL}_{2}({\mathbb{Z}})$ whose Fourier coefficients are the complex numbers $A_{m}$, $m\in{\mathbb{Z}}_{>0}$. Let $f\colon{\mathbb{H}}\to{\mathbb{C}}$ be the function
$f(y):=\frac{1}{2}g\left(-\frac{1}{y}\right)+\sum_{m>0}A_{m}\mathbf{e}_{1}\left(-\frac{m}{y}\right)\,,$ (99)
where $\mathbf{e}_{1}$ is defined in Eq. (91). Then, the Gevrey-1 asymptotic series $\tilde{f}(y)\in{\mathbb{C}}\llbracket y\rrbracket$ obtained expanding $f(y)$ for $y\to 0$ with $\Im(y)>0$ has a modular resurgent structure with simple poles at $\zeta_{m}=2\pi{\rm i}m$, $m\in{\mathbb{Z}}_{>0}$, and corresponding Stokes constants $-A_{m}$. In addition,
  • ${\cal S}^{\mathrm{med}}_{\frac{\pi}{2}}\tilde{f}(y)=f(y)$ for every $y\in{\mathbb{H}}$;
  • $f(y)$ is a weight $-\omega$ holomorphic quantum modular form for $\mathsf{SL}_{2}({\mathbb{Z}})$.
Proof.
Recall that the asymptotic expansion of the exponential function $\mathbf{e}_{1}(z)$ in Eq. (91) in the limit $z\to\infty$ is [42, (8.11.E2)]
$\mathbf{e}_{1}(z)=\frac{1}{2\pi{\rm i}}\sum_{k=0}^{\infty}\frac{(-1)^{k}k!}{(2\pi{\rm i}z)^{k+1}}\,.$ (100)
It follows that the formal power series $\tilde{f}$ reads
$\tilde{f}(y)=-\frac{1}{2\pi{\rm i}}\sum_{m>0}A_{m}\sum_{n=0}^{\infty}\frac{n!}{(2\pi{\rm i}m)^{n+1}}y^{n+1}=-\frac{1}{2\pi{\rm i}}\sum_{n=1}^{\infty}\sum_{m>0}\frac{A_{m}}{(2\pi{\rm i}m)^{n}}\Gamma(n)y^{n}\,.$ (101)
Hence, its perturbative coefficients are of the form in Eq. (24) with $\zeta_{m}=2\pi{\rm i}m$. By Proposition 3.1, the series $\tilde{f}$ has a simple resurgent structure with a tower of simple poles located at $\zeta_{m}$ for $m\in{\mathbb{Z}}_{>0}$ and corresponding Stokes constants $-A_{m}$. Moreover, since the $L$-function
$L_{g}(s):=\sum_{m>0}\frac{A_{m}}{m^{s}}$ (102)
admits a meromorphic continuation to the whole complex $s$-plane due to a well-known result by Hecke [43, Section 5], $\tilde{f}$ is modular resurgent.
Let us now compute the median resummation applying Eq. (15). Let $\Im(y)>0$ and fix a small angle $\theta>0$. We have that
${\cal S}^{\mathrm{med}}_{\frac{\pi}{2}}\tilde{f}(y)$ $=-\frac{1}{2\pi{\rm i}}\int_{0}^{{\rm i}{\rm e}^{{\rm i}\theta}\infty}{\rm e}^{-\zeta/y}\sum_{n=1}^{\infty}\sum_{m>0}\frac{A_{m}}{(2\pi{\rm i}m)^{n}}\zeta^{n-1}d\zeta+\frac{1}{2}\sum_{m>0}A_{m}{\rm e}^{-2\pi{\rm i}m/y}$
$=\frac{1}{2\pi{\rm i}}\sum_{m>0}A_{m}\,\int_{0}^{{\rm i}{\rm e}^{{\rm i}\theta}\infty}{\rm e}^{-\zeta/y}\frac{d\zeta}{\zeta-2\pi{\rm i}m}+\frac{1}{2}g\left(-\frac{1}{y}\right)$
$=\frac{1}{2\pi{\rm i}}\sum_{m>0}A_{m}\,\int_{0}^{\infty}{\rm e}^{-2\pi t}\frac{dt}{t-\frac{{\rm i}m}{y}}+\frac{1}{2}g\left(-\frac{1}{y}\right)\,,$
where we used the absolute convergence of the integrand in the second step. The desired result then follows from the definitions in Eqs. (91) and (99).
Finally, let us verify the quantum modularity property. For simplicity, we choose a pair of generators of $\mathsf{SL}_{2}({\mathbb{Z}})$ different from before, namely, $S$ and $\gamma:=TS$, where $T$ and $S$ are the standard generators in Eq. (81). Then, the corresponding cocycles are
$h_{S}[f](y)$ $=y^{\omega}\sum_{m>0}A_{m}\mathbf{e}_{1}(my)-\sum_{m>0}A_{m}\mathbf{e}_{1}\left(-\frac{m}{y}\right)\,,$ (103a)
$h_{\gamma}[f](y)$ $=y^{\omega}\sum_{m>0}A_{m}\mathbf{e}_{1}\left(-\frac{my}{y-1}\right)-\sum_{m>0}A_{m}\mathbf{e}_{1}\left(-\frac{m}{y}\right)\,,$ (103b)
respectively, where we implicitly used the modularity of the cusp form $g$. ∎
5 Conclusions
In this paper, we considered a particular class of resurgent asymptotic series that appear in numerous examples from the study of quantum invariants of knots and 3-manifolds, combinatorics, and topological string theory on toric CY threefolds. Motivated by the recent mathematical and physical interest, we introduced the notion of a modular resurgent series, whose Borel plane displays a single infinite tower of singularities, the secondary resurgent series are trivial, and the Stokes constants are coefficients of an $L$-function.
We proposed a new perspective on their modular resurgent structures by taking the point of view of the Stokes constants, their generating functions in the form of $q$-series, and their $L$-functions. Generalizing the strong-weak resurgent symmetry of the spectral trace of local ${\mathbb{P}}^{2}$ that we presented in the companion paper [20], we described the broader framework of modular resurgence. Pairs of modular resurgent structures are connected through a global exact symmetry revolving around the central role played by the functional equation satisfied by the $L$-functions whose coefficients are the Stokes constants. We illustrated the workings of modular resurgence via explicit computations in the concrete example of the generating functions of the Stokes constants for the weak and strong coupling asymptotic expansions of the spectral trace of local ${\mathbb{P}}^{2}$. Supported by disparate examples, we conjectured that the median resummation of modular resurgent series is effective and produces quantum modular forms. Finally, we proved that a large class of examples of modular resurgent series originating from the theories of Maass cusp forms and more general modular forms satisfies our conjectural statements.
Among the possible questions one could address, we highlight two research directions arising from the investigation performed in this paper. We want to broaden the supporting basis for our paradigm of modular resurgence. A bigger pool of examples might help us deepen our understanding of the critical role of the $L$-functions constructed from the Stokes constants and their functional equation. At the same time, a complete characterization of these ingredients might steer us towards proving Conjectures 1 and 2.
Moreover, our statements on the median resummation could be placed in a more general context. Indeed, as the first author started addressing in [33], when divergent formal power series arise from the asymptotic expansion of well-defined analytic functions, it is natural to ask which summability method allows us to reconstruct the original functions. It would be interesting to study the effectiveness of summability methods in a broad sense, following the spirit of our Conjectures 1 and 3 and Conjecture 1.1 in [1].
Acknowledgements
We thank Alba Grassi, Maxim Kontsevich, Marcos Mariño, Campbell Wheeler, and Don Zagier for many useful discussions. This work has been supported by the ERC-SyG project “Recursive and Exact New Quantum Theory” (ReNewQuantum), which received funding from the European Research Council (ERC) within the European Union’s Horizon 2020 research and innovation program under Grant No. 810573 and by the Swiss National Centre of Competence in Research SwissMAP (NCCR 51NF40-141869 The Mathematics of Physics).